### v3

# Distributed Optimization by Network Flows with Spatio-Temporal Compression

Abstract
Several data compressors have been proposed in distributed optimization frameworks of network systems to reduce communication overhead in large-scale applications. In this paper, we demonstrate that effective information compression may occur over time or space during sequences of node communications in distributed algorithms, leading to the concept of spatio-temporal compressors. This abstraction classifies existing compressors as spatio-temporal compressors, with their effectiveness described by constructive stability criteria from nonlinear system theory. Subsequently, we apply these spatio-temporal compressors to standard continuous-time consensus flows and distributed prime-dual flows, establishing conditions ensuring convergence. Additionally, we introduce a novel observer-based distributed primal-dual continuous flow integrated with spatio-temporal compressors, which provides broader convergence conditions. These continuous flows achieve exponential convergence to the global optimum when the objective function is strongly convex and can be discretized using Euler approximations. Finally, numerical simulations illustrate the versatility of the proposed spatio-temporal compressors and verify the convergence of algorithms.
keywords: Communication compression; Distributed optimization; Exponential convergence; Spatio-temporal compressors
1 Introduction
Distributed intelligent systems, such as drone swarms, smart grids, and cyber-physical systems, have been extensively researched across disciplines such as control, signal processing, and machine learning [2]-[5]. The mathematical representation of a distributed system involves a network connecting multiple agents, where each node symbolizes an individual agent, and the edges depict communication lines between these nodes. When distributed systems are required to implement tasks such as cluster optimization and collaborative control, it is required to compute distributively. In this process, each node stores localized information, communicates messages with connected nodes through the network, and collaboratively solves a mathematical problem [2]. This paper focuses on addressing distributed optimization problems, where each node possesses a function, aiming to identify solutions that collectively minimize the sum of all functions through communication across the network. Extensive effort has been devoted to developing distributed optimization algorithms, primarily rooted in the consensus algorithm. The goal of the consensus algorithm is to reach a consensus of the states across nodes. A combination of the consensus algorithm with the classical gradient descent method in optimization problems, coupled with stability tactics, results in the distributed (sub)gradient algorithm (DGA), achieving sublinear convergence under a strongly convex global objective function [6]-[8]. More algorithms have been introduced to address distributed optimization problems with faster rate requirements. For example, the distributed gradient tracking algorithm (DGT) incorporates an additional state to trace the gradient of the objective function [9, 10], akin to integral action [11]. Besides, for different formulations of distributed optimization problems, various Lagrangian functions have been proposed, giving rise to multiple algorithms based on the saddle point dynamic method. Examples include the Wang–Eila algorithm in [12] and the primal-dual algorithm in [13], distinct in communication states. These algorithms above can ensure linear convergence.
In practical implementation, the network bandwidth for communication in distributed systems is limited and numerous communication compression strategies have been developed to handle this issue. In [14]-[17], quantified communication is combined with DSG and DGT algorithms. Specifically, [14] introduced adaptive quantization and [15] applied random quantization, where the quantizer codebook changes when approaching the solution. In [17], the authors developed a dynamic encoding and decoding scheme for quantization. In addition to quantization, there are also several other types of compressors capable of reducing communication bits by synthesizing concepts from quantization, sparsity, and randomization [18]-[21]. Other researches aim to identify commonalities among specific compressors, thereby proposing a generalized definition of compressors [22]-[24]. The definition allows any function that satisfies these properties to be considered as a compressor and applied to algorithms. These compressors above exclusively focus on the spatial dimension, encompassing the information within transmitted messages. Notably, the compressor in [25] incorporates temporal dimensions, utilizing information across time. In addition to compressors, [26] has implemented an event-triggered method to reduce communication frequency, thereby alleviating the communication burden. This method could also be further developed into a form of compressor. Our goal is to propose such a definition, characterized by properties that simultaneously consider both temporal and spatial dimensions.
In addition, how to combine the compressors with distributed optimization algorithms has become a noteworthy area of study. This is because refining the application method can facilitate the successful integration of more general compressors and enhance the overall effectiveness of the algorithm. Beyond the direct application of compressors to the communication state, there exist other intriguing techniques, as direct application often poses challenges to stability [27, 28]. For instance, [14, 21] incorporate a weighted sum of the updated value and the original value into the original value, while [15, 29] compress the difference between iterations rather than the original value. In the work of [17, 30], the difference is scaled and then compressed, with the results communicated after a reverse reduction, further ensuring the convergence of the algorithm. [18, 31] adopts a difference application method based on filtering, and through additional equivalent transformations, ensures that only compressed values are exchanged. A compressor application method based on observers is also proposed in this paper.
This paper contributes in two aspects: proposing a general concept of compressors and exploring its application methods. Firstly, we introduce a general spatio-temporal (ST) compressor for communication compression in distributed optimization. This ST compressor leverages a constructive perspective on exponential stability from nonlinear systems, encompassing various communication compressors in the literature. We discuss two application methods: direct compression and observer-based compression, and establish convergence conditions for a class of distributed primal-dual algorithms with explicit convergence rates. Our results and analysis apply broadly to the ST communication compression framework, independent of the specific form of any particular compressor. Next, we propose the ST compressors in discrete-time cases, and the algorithms are established. We also compare these findings with error state application methods. Finally, we validate our conclusions through simulation experiments.
We begin with and focus on continuous-time systems in this paper, even though distributed optimization algorithms in real applications are mostly in discrete-time. This is because, first, the properties of our ST compressors are described by the induced systems and the converse Lyapunov theorem is used, with the continuous-time form being more intuitive for our analysis. Secondly, we would like to build our work as a natural extension of the excellent results on continuous-time distributed optimization algorithms in the control system community, e.g., [32]-[34]. In addition, we also offer a discrete version of the proposed continuous algorithm by Euler approximation, enabling our results to be applied in real situations. Our simulation experiments also illustrate the conclusion.
The paper is structured as follows. Section 2 formulates the distributed optimization problem of interest and proposes the spatio-temporal compressor of both the original form and the stronger form for message communication. In Section 3, we start from the distributed consensus to illustrate the conditions required when applying this compressor directly, as well as the successful combination with observer-based application methods. In Section 4, we respectively discuss the applicability of these two application methods to the primal-dual flow. In Section 5, we propose the ST compressors in discrete-time cases and discretize the continuous-time flows by the Euler method. Numerical simulations are presented to show the effectiveness of the proposed approaches in Section 6. Finally, a conclusion is made in Section 7. All proofs are collected in the Appendices. Compared to our preliminary version [1], in this paper, we introduce the concept of strong ST compressors and study continuous-time consensus flow. We also propose a compressor application method based on observers. Furthermore, we introduce the ST compressors in discrete-time cases, and propose two discrete algorithms based on Euler discretization.
[…]
2 Problem Formulation
In this section, we define the problem of interest.
2.1 Distributed Optimization
Consider a network of agents indexed by $\mathrm{V}=\{1,2,\dots,n\}$, where each agent $i\in\mathrm{V}$ holds an objective function $f_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}$. The agents aim to solve the following system-level optimization problem
$\mathrm{min}$ $\ \sum_{i=1}^{n}f_{i}(\mathbf{x}_{i})$ (1)
$\mathrm{s.t.}$ $\ \mathbf{x}_{i}=\mathbf{x}_{j},\quad\forall i,j\in\mathrm{V}.$
Particularly, each local objective function $f_{i}$ is assumed to fulfill the following requirements.
Assumption 1
The following properties are satisfied.
  • i)
    The global objective function $f(\mathbf{x}_{e}):=\sum_{i=1}^{n}f_{i}(\mathbf{x}_{e})$ is strongly convex, i.e., there exists $\mu>0$ such that $f(\mathbf{x}^{\prime}_{e})\geq f(\mathbf{x}_{e})+\nabla f(\mathbf{x}_{e})^{T}(\mathbf{x}^{\prime}_{e}-\mathbf{x}_{e})+\frac{\mu}{2}\|\mathbf{x}^{\prime}_{e}-\mathbf{x}_{e}\|^{2}$ for all $\mathbf{x}^{\prime}_{e},\mathbf{x}_{e}\in\mathbb{R}^{d}$.
  • ii)
    Each local gradient $\nabla f_{i}$ is globally Lipschitz continuous, i.e., there exists $L_{f}>0$ such that $\|\nabla f_{i}(\mathbf{x}_{e})-\nabla f_{i}(\mathbf{x}^{\prime}_{e})\|\leq L_{f}\|\mathbf{x}_{e}-\mathbf{x}^{\prime}_{e}\|$ for all $\mathbf{x}_{e},\mathbf{x}^{\prime}_{e}\in\mathbb{R}^{d}$. $\square$
If Assumption 1 holds, then the considered optimization problem (1) turns out a strongly convex optimization problem, allowing a unique solution $s^{\ast}\subseteq\mathbb{R}^{d}$ such that $\nabla f(s^{\ast})=0$ and $f(s^{\ast})=f^{\ast}$, where $f^{\ast}$ is the optimal value.
As each agent only has the information of its own objective function, to solve such a distributed optimization problem (1), a communication network is usually required for transmitting messages. Denote the communication graph $\mathrm{G=(V,E)}$, where $\mathrm{E}$ denotes the set of edges. Let $[a_{ij}]\in\mathrm{R}^{n\times n}$ denote the weight matrix, i.e., $a_{ij}>0$ if $(j,i)\in\mathrm{E}$ and $a_{ij}=0$ if $(j,i)\notin\mathrm{E}$. Then denote the Laplacian matrix of graph $\mathrm{G}$ by $\mathbf{L}$, satisfying $[\mathbf{L}]_{ij}=-a_{ij}$ for all $i\neq j$, and $[\mathbf{L}]_{ii}=\sum_{j=1}^{n}a_{ij}$ for all $i\in\mathrm{V}$. Denote the neighbor set of agent $i$ as $\mathrm{N}_{i}$, satisfying $j\in\mathrm{N}_{i}$ if and only if $[\mathbf{L}]_{ij}\neq 0$ for all $i,j\in\mathrm{V}$. For simplicity, we make the following assumption on the communication graph.
Assumption 2
The graph $\mathrm{G}$ is undirected and connected.
Assumption 2 indicates that the Laplacian matrix $\mathbf{L}$ is symmetric and positive semi-definite, with $[\mathbf{L}]_{ij}=[\mathbf{L}]_{ji}$, $\mathbf{L}\mathbf{1}_{n}=\mathbf{0}_{n}$ and its eigenvalues $\lambda_{i}$, $i\in\mathrm{V}$ in an ascending order satisfying $0=\lambda_{1}<\lambda_{2}\leq\dots\leq\lambda_{n}$ by [2]. We let $\mathbf{S}\in\mathbb{R}^{n\times(n-1)}$ be a matrix whose rows are eigenvectors corresponding to nonzero eigenvalues of $\mathbf{L}$, satisfying
$\mathbf{S}^{\top}\mathbf{1}_{n}=\mathbf{0}_{n-1}\,,\quad\mathbf{I}_{n}=\mathbf{S}\mathbf{S}^{\top}+\mathbf{1}_{n}\mathbf{1}_{n}^{\top}/n.$
Various distributed optimization algorithms have been developed in the literature [9]-[13] to compute the solution $s^{\ast}$ for (1). We mainly focus on the distributed primal-dual algorithm, which enables linear convergence and further generalizations to the case with constraints [35, 36]. A common distributed primal-dual flow for (1) takes the form [12, 13, 37]
$\dot{\mathbf{x}}_{i}$ $=-\sum^{n}_{j=1}\mathbf{L}_{ij}{\mathbf{x}}_{j}-\beta\mathbf{v}_{i}-\eta\nabla f_{i}(\mathbf{x}_{i}),$ (2)
$\dot{\mathbf{v}}_{i}$ $={\beta\sum^{n}_{j=1}\mathbf{L}_{ij}\mathbf{x}_{j}},$
where $\beta,\eta>0$ are parameters to be fixed and the initial condition $\sum_{i=1}^{n}\mathbf{v}_{i}(0)=\mathbf{0}_{d}$.
2.2 Spatio-Temporal Compressors
We propose the following spatio-temporal (ST) compressors for compressing node-to-node communications in distributed algorithms.
Definition 1 (ST compressor)
Given a uniformly linearly bounded mapping $\mathbf{C}:\mathbb{R}^{d}\times\mathbb{R}_{+}\rightarrow\mathbb{R}^{d}$, i.e., there exists a $L_{c}>0$ such that $\|\mathbf{C}(\mathbf{x}_{e},t)\|\leq L_{c}\|\mathbf{x}_{e}\|$ for all $\mathbf{x}_{e}\in\mathbb{R}^{d}$ and any $t\in\mathbb{R}_{+}$. The mapping $\mathbf{C}$ is said to be a ST compressor if the induced continuous-time non-autonomous system $\dot{\mathbf{x}}_{e}=-\mathbf{C}(\mathbf{x}_{e},t)$ is uniformly globally exponentially stable at the origin. $\square$
In the following, a stronger version of the spatio-temporal compressors, i.e., the strong spatio-temporal (SST) compressors, is introduced.
Definition 2 (SST compressor)
Given a uniformly globally Lipschitz mapping $\mathbf{C}:\mathbb{R}^{d}\times\mathbb{R}_{+}\rightarrow\mathbb{R}^{d}$, i.e., there exists a $L_{c}>0$ such that $\|\mathbf{C}(\mathbf{x}_{e},t)-\mathbf{C}(\mathbf{x}_{e}^{\prime},t)\|\leq L_{c}\|\mathbf{x}_{e}-\mathbf{x}_{e}^{\prime}\|$ for all $\mathbf{x}_{e},\mathbf{x}_{e}^{\prime}\in\mathbb{R}^{d}$ and any $t\in\mathbb{R}_{+}$. The mapping $\mathbf{C}$ is said to be a SST compressor if the induced continuous-time non-autonomous system $\dot{\mathbf{x}}_{e}=-k\mathbf{C}(\mathbf{x}_{e},t)$ is uniformly globally exponentially stable at the origin for any $k>0$. $\square$
It is clear that the mapping $\mathbf{C}$, satisfying either of both Definitions 1 and 2, needs to vanish at the origin, i.e., $\mathbf{C}(0,t)\equiv 0$ uniformly in $t$. This immediately shows that the uniformly globally Lipschitz property in Definition 2 implies the uniformly linearly bounded property. Besides, the exponential stability of the induced non-autonomous $\mathbf{x}_{e}$-systems in Definition 2 is more restrictive than in Definition 1. In view of both points, the notion of the SST compressor in Definition 2, is more restrictive than that of Definition 1.
We show that various compressors proposed in the literature can be categorized into ST or SST compressors in the following examples.
Example 1
The scalarization compressor $\mathbf{C}_{1}:\mathbb{R}^{d}\times\mathbb{R}_{+}\rightarrow\mathbb{R}^{d}$ satisfies $\mathbf{C}_{1}(\mathbf{x}_{e},t)=\bm{\psi}(t)\bm{\psi}(t)^{T}\mathbf{x}_{e}$, where the compression vector $\bm{\psi}:\mathbb{R}_{+}\rightarrow\mathbb{R}^{d}$ is uniformly bounded and persistently excited, i.e.,
[…]
for some constants $\alpha_{1},\alpha_{2},T_{1}>0$ (see [25]). $\square$
Example 2
The contraction compressor $\mathbf{C}_{2}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ satisfies
$\|\frac{\mathbf{C}_{2}(\mathbf{x}_{e})}{p}-\mathbf{x}_{e}\|^{2}\leq(1-\varphi)\|\mathbf{x}_{e}\|^{2},$ (4)
for some $\varphi\in(0,1]$ and $p>0$ (see [18, 21, 23], with the expectation operator removed). The following $\mathbf{C}_{2a}$ and $\mathbf{C}_{2b}$ are specific examples of $\mathbf{C}_{2}$ ($p=1$ and $\varphi=\frac{k}{d}$ of $\mathbf{C}_{2a}$, $p=\frac{d}{2}$ and $\varphi=\frac{1}{d^{2}}$ of $\mathbf{C}_{2b}$, $p=1$ and $\varphi=\frac{3}{4}$ of $\mathbf{C}_{2c}$):
  • 2a).
    Greedy (Top-k) sparsifier [38], which is given by $\mathbf{C}_{2a}(\mathbf{x}_{e})=\sum_{s=1}^{k}[\mathbf{x}_{e}]_{i_{s}}{\bf e}_{i_{s}}$ where $i_{1},\dots,i_{k}$ are the indices of largest $k$ coordinates in the absolute value of $\mathbf{x}_{e}$.
  • 2b).
    Standard uniform quantizer [18], which is given by $\mathbf{C}_{2b}(\mathbf{x}_{e})=\frac{\|\mathbf{x}_{e}\|_{\infty}}{2}\mathrm{sgn}(\mathbf{x}_{e}),$ where $\mathrm{sgn}(\cdot)$ denotes the element-wise sign.
  • 2c).
    Saturated quantizer, which is given by
    $[\mathbf{C}_{2c}(\mathbf{x}_{e})]_{i}=\left\{\begin{aligned} {}_{i},\quad\quad\ \ |[\mathbf{x}_{e}]_{i}|\leq\Delta.\\ \Delta\lfloor\frac{[\mathbf{x}_{e}]_{i}}{\Delta}\rfloor,\quad|[\mathbf{x}_{e}]_{i}|>\Delta.\\ \end{aligned}\right.$
[…]
Proposition 1
The following statements are true: a). $\mathbf{C}_{1}$ belongs to the SST compressor; b). $\mathbf{C}_{2}$ belongs to the ST compressor. $\square$
Remark 1
We stress that when the compressor $\mathbf{C}(\mathbf{x}_{e},t)$ is used, we do not mean to use $\mathbf{C}(\mathbf{x}_{e},t)$ to encode $\mathbf{x}_{e}$ for communication and then transmitting the whole vector of $\mathbf{C}$ directly. Instead, the $\mathbf{C}$ represents the communication information, whose transmission can be implemented requiring fewer bandwidths than directly transmitting $\mathbf{x}_{e}$ of $d$ dimensions, leading to the so-called communication compression. For example, if the scalarization compressor $\mathbf{C}_{1}$ is adopted, the actual communication message at each round is a scalar $\bm{\psi}(t)^{\top}\mathbf{x}_{e}(t)$ with each agent holding a common $\bm{\psi}(t)$, while for the standard uniform quantizer $\mathbf{C}_{2b}$, the actual communication message consists of a scalar $\|\mathbf{x}_{e}\|_{\infty}$ and a vector $\mathrm{sgn}(\mathbf{x}_{e})$. In view of this, with a bit of abuse of notation, we insist on saying the mapping $\mathbf{C}$ to be a compressor throughout the paper. $\square$
Remark 2
In contrast with the conventional compressors, e.g., the contraction compressor, the ST compressor exhibits two distinctive features. Firstly, it synthesizes information from both the time and space domains, broadening its applicability and expanding the design possibilities. Secondly, its key characteristic is elucidated through a non-autonomous system, which can simplify the design procedure while providing the flexibility to incorporate control-related tools into distributed optimization.
3 ST-Compressed Consensus Flows
Distributed consensus is a fundamental algorithm that acts as a subroutine in numerous distributed optimization problems. In this section, we investigate how to combine the ST compressors with the consensus algorithm, which motivates the subsequent developments of distributed optimization algorithms with ST compressors. Moreover, due to its convenience of analysis, we focus on the continuous-time distributed consensus, taking the form
[…]
It is clear that over the graph $\mathrm{G}$ under Assumption 2, each node state exponentially reaches consensus at the average ${\mathbf{x}}^{\ast}:=\frac{1}{n}\sum_{j=0}^{n}\mathbf{x}_{j}(0)$ [7].
3.1 Distributed Consensus with Direct Compression
An intuitive design of the compressed consensus algorithm is to directly replace the transmitted message $\mathbf{x}_{i}$ with the compressed one, leading to the following distributed consensus flow with direct compression (DC-DC) as
[…]
Then a natural question arises: given a ST compressor $\mathbf{C}$, whether or when the DC-DC flow (6) maintains linear convergence to the average. Before we answer this question, we make the following observation on the SST compressor. Given a SST compressor $\mathbf{C}$, it is clear that the system
$\dot{\mathbf{y}}_{e}=-\Lambda\mathcal{C}(\mathbf{y}_{e},t),$
where ${\mathbf{y}}_{e}\in\mathbb{R}^{(n-1)d}$, $\Lambda:=\mathrm{diag}(\lambda_{2},\dots,\lambda_{n})\otimes\mathbf{I}_{d}$ and $\mathcal{C}(\mathbf{y},t):=[\mathbf{C}^{T}(\mathbf{y}_{1},t),\dots,\mathbf{C}^{T}(\mathbf{y}_{n-1},t)]^{T}$, is uniformly globally exponentially stable at the zero equilibrium. By the converse Lyapunov Theorem for exponential stability [39, Theorem 4.14], this implies the existence of a Lyapunov function $V_{e}:\mathbb{R}^{(n-1)d}\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ such that
$c_{1}\|\mathbf{y}_{e}\|^{2}\leq V_{e}(\mathbf{y}_{e},t)$ $\leq c_{2}\|\mathbf{y}_{e}\|^{2},$ (7)
$\frac{\partial V_{e}}{\partial t}-\frac{\partial V_{e}}{\partial\mathbf{y}_{e}}\Lambda\mathcal{C}(\mathbf{y}_{e},t)$ $\leq-c_{3}\|\mathbf{y}_{e}\|^{2},$
$\ \|\frac{\partial V_{e}}{\partial\mathbf{y}_{e}}\|$ $\leq c_{4}\|\mathbf{y}_{e}\|,$
for some $c_{1},c_{2},c_{3},c_{4}>0$.
With this in mind, and defining $\mathbb{S}:=\mathbf{S}\otimes\mathbf{I}_{d}$ and $\mathcal{C}(\mathbf{x},t):=[\mathbf{C}^{T}(\mathbf{x}_{1},t),\dots,\mathbf{C}^{T}(\mathbf{x}_{n},t)]^{T}$, we are ready to propose the following theorem for flow (6), answering the question by showing that a condition on the communication network $\mathrm{G}$ and the SST compressor $\mathbf{C}$ is still required to maintain a linear convergence to the average.
Theorem 1
Let Assumption 2 hold, then for the DC-DC flow (6) with a SST compressor $\mathbf{C}$, if there holds
$\|\mathcal{C}(\mathbb{S}^{T}\mathbf{x},t)-\mathbb{S}^{T}\mathcal{C}(\mathbf{x},t)\|\leq\delta\|\mathbb{S}^{T}\mathbf{x}\|\,,\ \forall(\mathbf{x},t)\in\mathbb{R}^{nd}\times\mathbb{R}_{+}$ (8)
for $\delta<\frac{c_{3}}{c_{4}\lambda_{n}}$, then there holds $\|\mathbf{x}_{i}(t)-{\mathbf{x}}^{\ast}\|^{2}=\mathcal{O}(e^{-\gamma t})\,$ for $\gamma=\frac{c_{3}-c_{4}\delta\lambda_{n}}{c_{1}}$. $\square$
From the extra condition (8), it can be seen that the SST compressor $\mathbf{C}$ should satisfy some conditions about the network graph (see $\mathbb{S}$), which will ensure the linear convergence property of the DC-DC flow (6) in general. Moreover, by taking a linear form of SST compressor $\mathbf{C}(\mathbf{x}_{e},t)=M(t)\mathbf{x}_{e}$, e.g. the scalarization compressor $\mathbf{C}_{1}$, we note that the extra condition (8) reduces to
$\|[(\mathbf{I}_{n-1}\otimes M(t))\mathbb{S}^{T}-\mathbb{S}^{T}(\mathbf{I}_{n-1}\otimes M(t))]\mathbf{x}\|\leq\delta\|\mathbb{S}^{T}\mathbf{x}\|\,,$
cwhich holds for all $(\mathbf{x},t)\in\mathbb{R}^{nd}\times\mathbb{R}_{+}$, since $(\mathbf{I}_{n-1}\otimes M(t))\mathbb{S}^{T}-\mathbb{S}^{T}(\mathbf{I}_{n-1}\otimes M(t))=0$. This immediately implies that the linear SST compressor, e.g., the scalarization compressor $\mathbf{C}_{1}$, is applicable to the DC-DC flow (6) with no need of any extra condition. This flow with the scalarization compressor $\mathbf{C}_{1}$, as a special case of Theorem 1, is studied in [25].
3.2 Distributed Consensus with Observer-based Compression
In the previous section, it has been shown that the SST compressor can be directly applied, subject to an extra condition (8) which poses limitations on the range of feasible compressors and communication graphs. In this section, such limitations are removed by proposing a distributed compressed consensus, taking the form
[…]
where $\alpha>0$ is a gain parameter, and ${\hat{\mathbf{x}}}^{j}_{i}(0)={\hat{\mathbf{x}}}^{j^{\prime}}_{i}(0),\forall j,j^{\prime}\in\mathrm{N}_{i}$, $i\in\mathrm{V}$.
The proposed compressed consensus flow (9) is comprised of two sets of states for each agent $i$. The state $\mathbf{x}_{i}$ denotes the estimate of the consensus solution as in (5), while the states ${\hat{\mathbf{x}}}^{i}_{j}$ are introduced to each agent $i$ to estimate its neighboring solution state $\mathbf{x}_{j}$, $j\in\mathrm{N}_{i}$. To have a better view of this, let us first ignore the compressor and have ${\mathbf{x}}_{i,c}={\mathbf{x}}_{i}-{\hat{\mathbf{x}}}^{i}_{i}$ in (9). Then it is clear that the ${\hat{\mathbf{x}}}^{i}_{j}$ acts as an observer to estimate $\mathbf{x}_{j}$. Our proposed compression strategy is thus established by compressing the observation error ${\mathbf{x}}_{i}-{\hat{\mathbf{x}}}^{i}_{i}$ between the solution state ${\mathbf{x}}_{i}$ and the corresponding observer state ${\hat{\mathbf{x}}}^{i}_{i}$ of agent $i$, as the message for communication with neighbors.
We are ready to propose the following theorem for flow (9).
Theorem 2
Let Assumption 2 hold, then for the DC-OC flow (9) with a ST compressor $\mathbf{C}$, there exists $\alpha^{\ast}=\mathrm{min}\{\frac{2c_{3}}{9\lambda_{n}c_{4}\sqrt{n}},\frac{2c_{3}}{3\lambda_{n}}\}$ such that for all $\alpha\leq\alpha^{\ast}$, there holds $\|\mathbf{x}_{i}(t)-{\mathbf{x}}^{\ast}\|^{2}=\mathcal{O}(e^{-\gamma t})\,,$ for some $\gamma=\mathrm{min}\{\frac{\alpha\lambda_{2}}{2},\frac{c_{3}}{3c_{1}}\}$. (see Appendix .3 for the parameters $c_{1}$,$c_{3}$,$c_{4}$.) $\square$
A rigorous proof of Theorem 2 is presented in Appendix C. Intuitively, from the perspective of control systems, we stress that the corresponding system (9) can be regarded as an interconnection of two subsystems: $\mathbf{x}_{i}$-subsystem and ${\hat{\mathbf{x}}}^{i}_{j}$-subsystem, with $\alpha$ a low gain that is tuned to be small such that the supply functions of the two interconnected subsystems satisfy some small-gain conditions for closed-loop exponential stability [39, Theorem 5.6].
4 ST-Compressed primal-dual Flows
4.1 Direct Compression
In this section, we aim to present a distributed compressed primal-dual flow for problem (1) by applying the SST compressor to directly compress the communication message in the conventional primal-dual algorithms, as in Section 3.1.
The proposed distributed primal-dual flow with direct compression takes the form
$\dot{\mathbf{x}}_{i}$ $=-\sum^{n}_{j=1}\mathbf{L}_{ij}\mathbf{C}(\mathbf{x}_{j},t)-\beta\mathbf{v}_{i}-\eta\nabla f_{i}(\mathbf{x}_{i}),$ (10)
$\dot{\mathbf{v}}_{i}$ $={\beta\sum^{n}_{j=1}\mathbf{L}_{ij}\mathbf{C}(\mathbf{x}_{j},t)},$
where the initial condition $\sum_{i=1}^{n}\mathbf{v}_{i}(0)=\mathbf{0}_{d}$.
We are ready to propose the following theorem for flow (10).
Theorem 3
Let Assumptions 1 and 2 hold, and $\mathbf{C}$ be a SST compressor, which also satisfies (8) with some $\delta>0$. Then there exist some $\beta,\eta>0$ such that $\mathbf{x}_{i}(t)$ generated by flow (10) converges to the optimal solution $s^{\ast}$ exponentially (see Appendix .4 for the parameters $\delta$,$\beta$,$\eta$). $\square$
This theorem demonstrates the effectiveness of direct compression. When the conditions are satisfied, the direct application of the compressor can also ensure exponential convergence, without the need to resort to complex application methods.
4.2 Observer-based Compression
In this section, we propose another compressed distributed primal-dual flow by applying the ST compressor to compress the communication message in the conventional primal-dual algorithms, based on distributed observer-based compressed consensus (9) in Section 3.2.
The proposed distributed primal-dual flow in continuous-time form with observer-based compression takes the form
$\dot{\mathbf{x}}_{i}$ $=-\alpha\sum^{n}_{j=1}\mathbf{L}_{ij}{\hat{\mathbf{x}}}^{i}_{j}-\beta\mathbf{v}_{i}-\eta\nabla f_{i}(\mathbf{x}_{i}),$ (11)
$\dot{\mathbf{v}}_{i}$ $={\beta\sum^{n}_{j=1}\mathbf{L}_{ij}{\hat{\mathbf{x}}}^{i}_{j}},$
$\dot{\hat{\mathbf{x}}}^{i}_{j}$ $={\mathbf{x}}_{j,c},\quad j\in\mathrm{N}_{i}$
${\mathbf{x}}_{i,c}$ $=\mathbf{C}({\mathbf{x}}_{i}-{\hat{\mathbf{x}}}^{i}_{i},t),$
where the initial condition is $\sum_{i=1}^{n}\mathbf{v}_{i}(0)=\mathbf{0}_{d}$ and for every $i$, ${\hat{\mathbf{x}}}^{j}_{i}(0)={\hat{\mathbf{x}}}^{j^{\prime}}_{i}(0),\forall j,j^{\prime}\in\mathrm{V}$.
We are ready to propose the following theorem for Flow (11).
Theorem 4
Let Assumptions 1 and 2 hold, and $\mathbf{C}$ be a ST compressor. Then there exist $\alpha,\beta,\eta>0$ such that $\mathbf{x}_{i}(t)$ generated by flow (11) converges to the optimal solution $s^{\ast}$ exponentially (see Appendix .5 for the parameters $\alpha$,$\beta$,$\eta$). $\square$
5 Discrete Implementations
5.1 ST Compressors in Discrete Time
Based on the ST compressors in Definitions 1 and 2, we propose the following ST compressors in discrete time.
Definition 3 (ST compressor in discrete time)
Given a uniformly linearly bounded mapping $\mathbf{C}:\mathbb{R}^{d}\times\mathbb{R}_{+}\rightarrow\mathbb{R}^{d}$. The mapping $\mathbf{C}$ is said to be a ST compressor in discrete time if the induced discrete-time non-autonomous system $\mathbf{x}_{e}(t+1)=\mathbf{x}_{e}(t)-\kappa_{0}\mathbf{C}(\mathbf{x}_{e}(t),t)$ is uniformly globally linearly stable at the origin for some stepsize $\kappa_{0}>0$. $\square$
Definition 4 (SST compressor in discrete time)
Given a uniformly globally Lipschitz mapping $\mathbf{C}:\mathbb{R}^{d}\times\mathbb{R}_{+}\rightarrow\mathbb{R}^{d}$, The mapping $\mathbf{C}$ is said to be a SST compressor in discrete time if there exists $\kappa_{0}^{\ast}>0$ such that the induced discrete-time non-autonomous system $\mathbf{x}_{e}(t+1)=\mathbf{x}_{e}(t)-\kappa_{0}\mathbf{C}(\mathbf{x}_{e}(t),t)$ is uniformly globally linearly stable at the origin for any stepsize $\kappa_{0}\in(0,\kappa_{0}^{\ast})$. $\square$
In the discrete-time case, it should be noticed that the condition (3) of the scalarization compressor becomes $\alpha_{2}\mathbf{I}_{d}\geq\sum_{s=t}^{t+T_{1}-1}\bm{\psi}(s)\bm{\psi}^{\top}(s)\geq\alpha_{1}\mathbf{I}_{d}\,,\ \forall t\geq 0\,.$ A specific example of discrete-time cases of $\mathbf{C}_{1}$, denoted by $\mathbf{C}_{1a}$, can be derived by letting $\bm{\psi}(t)=\mathbf{e}_{i}$ with $i=1+(t\ \mathrm{mod}\ d)$ for $t\in\mathbb{N}$. Moreover, a similar proposition still holds.
Proposition 2
The following statements are true: a). $\mathbf{C}_{1}$ belongs to the SST compressor in discrete time; b). $\mathbf{C}_{2}$ belongs to the ST compressor in discrete time. $\square$
5.2 Euler Discretization of Compressed primal-dual Flows
In practice, algorithms are always implemented in a discrete-time form. In the following, we discretize the flow (10) based on the Euler method, and derive Algorithm 1.
Algorithm 1 Distributed primal-dual flow with direct compression (DPD-DC)
Initialization: $\kappa,\kappa_{0},\beta,\eta>0$, $\sum_{i=1}^{n}\mathbf{v}_{i}(0)=\mathbf{0}_{d}$.
for $t\in\mathbb{N}_{+}$, every node $i$ do
     Transmit $\mathbf{C}(\mathbf{x}_{j}(t),t)$ to neighbors
     Update:
     $\mathbf{x}_{i}(t+1)=\mathbf{x}_{i}(t)-\kappa_{0}\sum^{n}_{j=1}\mathbf{L}_{ij}\mathbf{C}(\mathbf{x}_{j}(t),t)-\kappa[\beta\mathbf{v}_{i}(t)\newline ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\eta\nabla f_{i}(\mathbf{x}_{i}(t))],$
     $\mathbf{v}_{i}(t+1)=\mathbf{v}_{i}(t)+\kappa_{0}{\beta\sum^{n}_{j=1}\mathbf{L}_{ij}\mathbf{C}(\mathbf{x}_{j}(t),t)}.$
end for
Theorem 5
Let Assumptions 1 and 2 hold, and $\mathbf{C}$ be a SST compressor in discrete time, which also satisfies (8) with some $\delta>0$. Then there exist some $\kappa,\kappa_{0},\beta,\eta>0$ such that $\mathbf{x}_{i}(t)$ generated by DPD-DC converges to the optimal solution $s^{\ast}$ linearly. $\square$
Next, we discretize Flow (11) based on the Euler approximation method, yielding Algorithm 2.
Algorithm 2 Distributed primal-dual flow with observer-based compression (DPD-OC)
Initialization: $\kappa,\kappa_{0},\beta,\eta>0$, $\sum_{i=1}^{n}\mathbf{v}_{i}(0)=\mathbf{0}_{d}$. Every node $i$ generates the observer of other nodes $\mathbf{x}_{j}^{i}(0)$, $j\in\mathrm{V}$. The initial value of different observers of node $i$ in each node $j$ is the same, i.e., ${\hat{\mathbf{x}}}^{j}_{i}(0)={\hat{\mathbf{x}}}^{j^{\prime}}_{i}(0),\forall j,j^{\prime}\in\mathrm{N}_{i}$, $i\in\mathrm{V}$.
for $t\in\mathbb{N}_{+}$, every node $i$ do
     Transmit $\mathbf{x}_{i,c}(t)=\mathbf{C}(\mathbf{x}_{i}(t)-{\hat{\mathbf{x}}}^{i}_{i}(t),t)$ to neighbors
     for every observer of node $j$ do
         ${\hat{\mathbf{x}}}^{i}_{j}(t+1)={\hat{\mathbf{x}}}^{i}_{j}(t)+\kappa_{0}\mathbf{x}_{j,c}(t)$
     end for
     Update:
     $\mathbf{x}_{i}(t+1)=\mathbf{x}_{i}(t)-\kappa[\sum_{j=1}^{n}\mathbf{L}_{ij}{\hat{\mathbf{x}}}^{i}_{j}(t)+\beta\mathbf{v}_{i}(t)\newline ~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\eta\nabla f_{i}(\mathbf{x}_{i}(t))],$
     $\mathbf{v}_{i}(t+1)=\mathbf{v}_{i}(t)+\kappa{\beta\sum_{j=1}^{n}\mathbf{L}_{ij}{\hat{\mathbf{x}}}^{i}_{j}(t)}.$
end for
Theorem 6
Let Assumptions 1 and 2 hold, and $\mathbf{C}$ be a ST compressor in discrete time with some $\kappa_{0}>0$. Then there exist some $\kappa,\beta,\eta>0$ such that $\mathbf{x}_{i}(t)$ generated by DPD-OC converges to the optimal solution $s^{\ast}$ linearly. $\square$
5.3 Compared with Error State Compression
Let us discuss the combination of our ST compressor with a commonly used method [18, 26, 31]. This method introduces a distributed filter and a distributed integrator and compresses the state errors. The proposed distributed primal-dual flow with error state compression (DPD-ESC) takes the form
${\bm{\sigma}}_{i}(t+1)$ $=\bm{\sigma}_{i}(t)+\kappa_{0}\mathbf{q}_{i}(t),$
${\mathbf{z}}_{i}(t+1)$ $=\mathbf{z}_{i}(t)+\kappa_{0}[\mathbf{q}_{i}(t)-\sum^{n}_{j=1}\mathbf{L}_{ij}\mathbf{q}_{j}(t)],$
${\mathbf{x}}_{i}(t+1)$ $=\mathbf{x}_{i}(t)-\kappa[(\bm{\sigma}_{i}(t)-\mathbf{z}_{i}(t)+\sum^{n}_{j=1}\mathbf{L}_{ij}{\mathbf{q}}_{j}(t))$
$\quad+\beta\mathbf{v}_{i}(t)+\eta\nabla f_{i}(\mathbf{x}_{i}(t))],$
${\mathbf{v}}_{i}(t+1)$ $=\mathbf{v}_{i}(t)+\kappa{\beta(\bm{\sigma}_{i}(t)-\mathbf{z}_{i}(t)+\sum^{n}_{j=1}\mathbf{L}_{ij}{\mathbf{q}}_{j}(t))},$
$\mathbf{q}_{i}(t)$ $=\mathbf{C}(\mathbf{x}_{i}(t)-\bm{\sigma}_{i}(t),t).$
The ST compressors can also be applied to this algorithm and similar theorems can be proved. For readers of interest, the validity of the continuous-time form of the algorithm above is proved in our preliminary work [1].
5.4 Stochastic ST Compressors and Algorithms
It should be noticed that many literature on compressor assumption take into account the presence of randomness. Therefore, we extend the ST compressor to randomness and research its effectiveness in applications. In this section, we study the randomization of the ST compressor and the application of DPD-OC as a example.
Introduce randomness to Definition 1, we obtain the definition of Stochastic Spatio-Temporal (StST) Compressor, with focus on discrete time.
Definition 5 (StST Compressor)
Given a linearly mean-square bounded mapping $\mathbf{C}:\mathbb{R}^{d}\times\mathbb{R}_{+}\rightarrow\mathbb{R}^{d}$, i.e., there exists a $L_{c}>0$ such that $\mathbb{E}\|\mathbf{C}(\mathbf{x}_{e},t)\|^{2}\leq L^{2}_{c}\|\mathbf{x}_{e}\|^{2}$ for all $\mathbf{x}_{e}\in\mathbb{R}^{d}$ and any $t\in\mathbb{R}_{+}$. Then, $\mathbf{C}$ is said to be a StST compressor, if the induced non-autonomous system $\mathbf{x}_{e}(t+1)=\mathbf{x}_{e}(t)-\kappa_{0}\mathbf{C}(\mathbf{x}_{e},t)$ is uniformly globally exponentially stable at the origin in the mean-square sense [^1] See [40] for the definition of mean square convergence, for some stepsize $\kappa_{0}>0$. $\square$
The ST compressor is a special case of the StST compressor. Moreover, some compressor assumptions in literature belongs to the StST compressor.
Example 3
The stochastic contractive compressor $\mathbf{C}_{3}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ satisfies
$\mathbb{E}\|\frac{\mathbf{C}_{3}(\mathbf{x}_{e})}{p}-\mathbf{x}_{e}\|^{2}\leq(1-\varphi)\|\mathbf{x}_{e}\|^{2}$ (12)
for some $\varphi\in(0,1]$ and $p>0$. By [18], the followings are specific examples of $\mathbf{C}_{3}$:
  • 3a).
    Unbiased $l$-bits quantizer [41]
    $\mathbf{C}_{3a}(\mathbf{x}_{e})=\frac{\|\mathbf{x}_{e}\|_{\infty}}{2^{l-1}}{\rm sign}(\mathbf{x}_{e})*\lfloor\frac{2^{l-1}|\mathbf{x}_{e}|}{\|\mathbf{x}_{e}\|_{\infty}}+\overline{\omega}\rfloor,$
    where $\overline{\omega}$ is a random perturbation vector uniformly sampled from $[0,1]^{d}$.
Proposition 3
Compressor $\mathbf{C}_{3}$ belongs to the StST compressor. $\square$
The proof of Proposition 2 is similar to that of Proposition b). and is omitted for simplicity.
We apply the StST compressor to DPD-OC and propose the following theorem for DPD-OC.
Theorem 7
Let Assumption 1 and 2 hold, and $\mathbf{C}$ be a StST compressor with some $\kappa_{0}>0$. Then for $\kappa,\beta,\eta>0$, the mean square of $\mathbf{x}_{i}(t)$ in the DPD-OC converges to the optimal solution $s^{\ast}$ linearly. $\square$
6 Numerical Simulations
6.1 Verification of ST Compressors
In this section, we verify that the compressors mentioned in this paper, $\mathbf{C}_{1a}$, $\mathbf{C}_{2a}$, $\mathbf{C}_{2b}$, $\mathbf{C}_{2c}$, $\mathbf{C}_{3a}$, satisfy the core property, i.e., the exponential stability of induced systems, of the ST compressor. Specifically, we let $d=5$, $k=2$ for $\mathbf{C}_{2a}$, $\Delta=1$ for $\mathbf{C}_{2c}$ and $l=4$ for $\mathbf{C}_{3a}$.
The figures respectively demonstrate the exponential convergence system $\dot{\mathbf{x}}_{e}=-\mathbf{C}(\mathbf{x}_{e},t)$ with different compressors, validating our conclusions. Other properties of the ST compressor can also be easily verified through theoretical analysis. Additionally, the satisfaction of these compressors with the conditions of the ST compressor in discrete-time form can also be verified through simulation, but it is omitted here.
6.2 Simulations under Different application Methods
In this section, we consider a network of $n$ nodes over a circle communication graph and the dimension of the local state is $d$, where each edge is assigned with the same unit weight and each node holds a local function $f_{i}(\mathbf{x}_{i})=\frac{1}{2}\|\mathbf{H}^{T}_{i}\mathbf{x}_{i}-b_{i}\|^{2}$ with some randomly generated $\mathbf{H}_{i}\in\mathbb{R}^{d}$ and $b_{i}\in\mathbb{R}$. Assume that the linear equation $\mathbf{H}x=\mathbf{b}$ has a unique solution $s^{\ast}$, where $\mathbf{H}=[\mathbf{H}_{1}\,\dots\,\mathbf{H}_{n}]^{\top}\in\mathbb{R}^{n\times d}$ and $\mathbf{b}=[b_{1}\,\dots\,b_{n}]^{\top}\in\mathbb{R}^{d}$, then we can conclude that the functions $f_{i}(\mathbf{x}_{i})$ satisfy Assumption 1 with $\mu>0$ and optimal solution $s^{\ast}$. Specifically, we let $n=10$, $d=5$ and $s^{\ast}=[2,-4,-4,2,-3]^{T}$. Next, we will apply different application methods to the compressors and compare their effects.
We use the scalar compressor $\mathbf{C}_{1a}$ and greedy sparsifier compressor $\mathbf{C}_{2a}$ as examples. In this application, we integrate DPD-DC , DPD-OC , DPD-ESC with $\mathbf{C}_{1a}$, and integrate DPD-OC , DPD-ESC , with $\mathbf{C}_{2a}$. Specifically, we let the parameters $\kappa_{0}=1$, $\kappa=0.002$, $\alpha=1$, $\beta=10$, $\eta=5$. The plot illustrates the sum of squared distances from the current $\mathbf{x}_{i}(t)$ to $s^{\ast}$, denoted as $\sum_{i=1}^{n}\|\mathbf{x}_{i}(t)-s^{\ast}\|^{2}$ over time. We also simulated the primal-dual algorithm in discrete time without a compressor under the same parameters for comparison. Notably, the algorithms exhibit linear convergence to the optimal solution, verifying the theorems.
6.3 Simulations with Convex Objective Functions
Next, we discuss the case where the objective function is convex but not strongly convex, while other settings are the same as that in the previous section. We take the objective function from [42] as
$f_{i}(\mathbf{x})=\sum_{j=1}^{d-1}\left[100([\mathbf{x}]_{j+1}-[\mathbf{x}]_{j}^{2})^{2}+([\mathbf{x}]_{j}-1)^{2}\right],$
whose optimal solution is $s^{\ast}=\mathbf{1}_{d}$. The simulation results of DPD-DC , DPD-OC , DPD-ESC with $\mathbf{C}_{1a}$, as examples, are shown. From the figure, we can see that the above algorithm can achieve asymptotic convergence to the optimal solution for convex but not strongly convex functions.
7 Conclusion
In this paper, we have introduced a type of spatio-temporal compressors that integrates both spatial and temporal characteristics, and effectively compresses information by leveraging information from both the time and space domains. This type of compressors has covered several compressors in the literature. Our proposed compressor has been implemented in the primal-dual algorithm by direct compression method and observer-based compression method. In the future, we will investigate a broader spectrum of compressor types and algorithms tailored to the characteristics of the compressors. We will also extend the application of the ST compressor to more classical distributed optimization algorithms, testing its universality across different algorithms.
.1 Proof of Proposition 1
Proof of a). The proof of the statement in Definition 2 is obvious by recalling [43] that the system $\dot{\mathbf{x}}_{e}=-k\bm{\psi}(t)\bm{\psi}(t)^{T}\mathbf{x}_{e}$ is globally exponentially stable at the zero equilibrium for any $k>0$. The proof of the linear boundedness property can be shown by noting that $\bm{\psi}(t)$ is uniformly bounded.
Proof of b). We proceed to show the compressor $\mathbf{C}_{2}(\mathbf{x}_{e})$ satisfies properties in Definition 1. Note that the contraction compressor (4) is equivalent to
$\|{\mathbf{C}_{2}(\mathbf{x}_{e})}/{p}\|^{2}-2{\mathbf{x}_{e}^{T}\mathbf{C}_{2}(\mathbf{x}_{e})}/{p}\leq-\varphi\|\mathbf{x}_{e}\|^{2}\,.$ (13)
First, we prove that $\mathbf{C}_{2}$ is the ST compressor by proving the system $\dot{\mathbf{x}}_{e}=-\mathbf{C}_{2}(\mathbf{x}_{e},t)$ is exponentially stable at the zero equilibrium. By choosing the Lyapunov function $V_{e}(\mathbf{x}_{e})={\|\mathbf{x}_{e}\|^{2}}/{p}$ and using (13), we have $\dot{V}_{e}=-2\frac{\mathbf{x}_{e}^{T}\mathbf{C}_{2}(\mathbf{x}_{e})}{p}\leq-\varphi\|\mathbf{x}_{e}\|^{2}.$ Thus $\mathbf{x}_{e}$-system is globally exponentially stable at the zero equilibrium with $\varphi>0$. Then, by (13) and using the Young’s inequality, we have
$\|{\mathbf{C}_{2}(\mathbf{x}_{e})}/{p}\|^{2}$ $\leq\frac{1}{2}\|{\mathbf{C}_{2}(\mathbf{x}_{e})}/{p}\|^{2}-(\varphi-2)\|\mathbf{x}_{e}\|^{2}$ (14)
$\Rightarrow\|{\mathbf{C}_{2}(\mathbf{x}_{e})}\|$ $\leq p\sqrt{{2(2-\varphi)}}\|\mathbf{x}_{e}\|\leq 2p\|\mathbf{x}_{e}\|,$
where the last inequality is obtained by $\varphi\in(0,1]$. Thus the linearly boundedness property is proved with $L_{c}=2p>0$. This completes the proof.
.2 Proof of Theorem 1
Flow (6) can be written in a tight form as
$\dot{\mathbf{x}}$ $=-\mathcal{L}\mathbf{x}_{c},$ (15)
$\mathbf{x}_{c}$ $=\mathcal{C}(\mathbf{x},t),$
where $\mathbf{x}:=[\mathbf{x}_{1}^{T},\dots,\mathbf{x}_{n}^{T}]^{T}$, $\mathbf{x}_{c}:=[\mathbf{x}_{1,c}^{T},\dots,\mathbf{x}_{n,c}^{T}]^{T}$ and $\mathcal{L}:=\mathbf{L}\otimes\mathbf{I}_{d}$. We decompose $\mathbf{x}$ by defining $\mathbf{x}_{\perp}:=\mathbb{S}^{T}\mathbf{x}=[\mathbf{x}^{T}_{\perp,1},\dots,\mathbf{x}^{T}_{\perp,n-1}]^{T}$ and $\mathbf{x}_{\parallel}:=\mathbb{I}^{T}\mathbf{x}$, where $\mathbb{I}:=\frac{1}{\sqrt{n}}\mathbf{1}_{n}\otimes\mathbf{I}_{d}$. We can obtain that $\dot{\mathbf{x}}_{\parallel}=\mathbf{0}_{d}$ in (15) by the fact
$\mathbb{I}^{T}\mathcal{L}=\mathbf{0}\quad\mathcal{L}\mathbb{I}=\mathbf{0}.$ (16)
Then with the fact
$\mathbb{S}\mathbb{S}^{T}+\mathbb{I}\ \mathbb{I}^{T}=\mathbf{I}_{nd},$ (17)
it can be noticed that we can prove $\mathbf{x}_{i}(t)$ converges to average consensus by proving $\mathbf{x}_{\perp}(t)$ converges to zero equilibrium.
We can obtain the derivation of $\mathbf{x}_{\perp}$ along with time
$\dot{\mathbf{x}}_{\perp}=-\mathbb{S}^{T}\mathcal{L}\mathbf{\mathcal{C}}(\mathbf{x},t).$ (18)
Define the Lypanuov function of (18) $V(\mathbf{x}_{\perp},t):=V_{e}(\mathbf{x}_{\perp},t)$, which is defined in (7), then we have
$\dot{V}$ $=\frac{\partial V}{\partial t}-\frac{\partial V}{\partial\mathbf{x}_{\perp}}\Lambda\mathcal{C}({\mathbf{x}}_{\perp},t)+\frac{\partial V}{\partial\mathbf{x}_{\perp}}[\Lambda\mathcal{C}({\mathbf{x}}_{\perp},t)-\mathbb{S}^{T}\mathcal{L}\mathbf{\mathcal{C}}(\mathbf{x},t)]$ (19)
$\leq-(c_{3}-c_{4}\delta\lambda_{n})\|\mathbf{x}_{\perp}\|^{2},$
where the inequality is obtained by the fact
$\|\Lambda\mathcal{C}(\mathbf{y}_{e},t)-\mathbb{S}^{T}\mathcal{L}\mathbf{\mathcal{C}}(\mathbf{x},t)\|$ $\leq\lambda_{n}\|\mathcal{C}({\mathbf{x}}_{\perp},t)-\mathbb{S}^{T}\mathcal{C}(\mathbf{x},t)\|$
$\leq\delta\lambda_{n}\|\mathbf{x}_{\perp}\|.$
If $\delta\leq\frac{c_{3}}{c_{4}\lambda_{n}}$, $\dot{V}$ is negative definite. With (7), we have $\dot{V}\leq-\frac{c_{3}-c_{4}\delta\lambda_{n}}{c_{1}}V,$ then $\|\mathbf{x}_{\perp}(t)\|^{2}=\mathcal{O}(e^{-\gamma t})$, where $\gamma=\frac{c_{3}-c_{4}\delta\lambda_{n}}{c_{1}}$. Thus the theorem is proved.
.3 Proof of Theorem 2
From Flow (9) and its initial condition, we can obtain that for every $i\in\mathrm{V}$, ${\hat{\mathbf{x}}}^{i}_{j}(0)={\hat{\mathbf{x}}}^{j^{\prime}}_{i}(0),\forall j,j^{\prime}\in\mathrm{V}$, i.e., the stored value of $\mathbf{x}_{i}$ is the same in every node. Thus the stored value of each node can be written as $\mathbf{x}_{c}:=[\mathbf{x}_{1,c}^{T},\dots,\mathbf{x}_{n,c}^{T}]^{T}$. Then Flow (9) can be written in a tight form as
$\dot{\mathbf{x}}$ $=-\alpha\mathcal{L}\mathbf{x}_{c},$ (20)
$\dot{\mathbf{x}}_{c}$ $=\mathcal{C}(\mathbf{x}-\mathbf{x}_{c},t),$
where $\mathcal{C}(\mathbf{x}-\mathbf{x}_{c},t):=[\mathbf{C}^{T}(\mathbf{x}_{1}-\mathbf{x}_{1,c},t),\dots,\mathbf{C}^{T}(\mathbf{x}_{n}-\mathbf{x}_{n,c},t)]^{T}.$ Similarly, we decompose $\mathbf{x}$ by defining $\mathbf{x}_{\perp}:=\mathbb{S}^{T}\mathbf{x}$ and $\mathbf{x}_{\parallel}:=\mathbb{I}^{T}\mathbf{x}$. Still $\dot{\mathbf{x}}_{\parallel}=\mathbf{0}_{d}$ and we will prove the convergence of $\mathbf{x}_{\perp}(t)$. By (20), we have
$\dot{\mathbf{x}}_{\perp}$ $=-\alpha\mathbb{S}^{T}\mathcal{L}\mathbf{x}_{c},$ (21)
$\dot{\mathbf{x}}_{c}$ $=\mathcal{C}(\mathbf{x}-\mathbf{x}_{c},t).$
We choose $V_{1}(\mathbf{x}_{\perp}):=\frac{1}{2}\|\mathbf{x}_{\perp}\|^{2}$, then we have
$\dot{V}_{1}\leq\frac{\alpha}{2}(-\lambda_{2}\|\mathbf{x}_{\perp}\|^{2}+\lambda_{n}\|\mathbf{x}-\mathbf{x}_{c}\|^{2}).$ (22)
As $\dot{\mathbf{x}}_{e}=-\mathbf{C}(\mathbf{x}_{e},t)$ is exponentially convergent at the zero equilibrium, where $\mathbf{x}_{e}\in\mathbb{R}^{d}$, then there exists a Lyapunov function $V_{e}(\mathbf{x}_{e},t):\mathbb{R}^{d}\times\mathbb{R}_{+}\rightarrow\mathbb{R}$ which satisfies
$c_{1}\|\mathbf{x}_{e}\|^{2}\leq V_{e}(\mathbf{x}_{e},t)$ $\leq c_{2}\|\mathbf{x}_{e}\|^{2},$ (23)
$\frac{\partial V_{e}}{\partial t}-\frac{\partial V_{e}}{\partial\mathbf{x}_{e}}\mathbf{C}(\mathbf{x}_{e},t)$ $\leq-c_{3}\|\mathbf{x}_{e}\|^{2},$
$\|\frac{\partial V_{e}}{\partial\mathbf{x}_{e}}\|$ $\leq c_{4}\|\mathbf{x}_{e}\|,$
for some $c_{1},c_{2},c_{3},c_{4}>0$.
We choose $V_{2}(\mathbf{x}-\mathbf{x}_{c},t):=\sum_{i=1}^{n}V_{e}(\mathbf{x}_{i}-\mathbf{x}_{i,c},t)$, then we have
$\dot{V}_{2}$ $\leq-c_{3}\|\mathbf{x}-\mathbf{x}_{c}\|^{2}+\frac{\alpha}{2}\lambda_{n}c_{4}\sqrt{n}\|\mathbf{x}-\mathbf{x}_{c}\|^{2}+\frac{\alpha}{2}c_{4}\sqrt{n}\|\mathbf{x}_{c}\|_{L}^{2}$ (24)
$\leq-(c_{3}-\frac{3\alpha}{2}\lambda_{n}c_{4})\|\mathbf{x}-\mathbf{x}_{c}\|^{2}+\alpha\lambda_{n}c_{4}\sqrt{n}\|\mathbf{x}_{\perp}\|^{2},$
where the first inequality is obtained by (23) and the second inequality is obtained by the fact
$\|\mathbf{x}_{c}\|_{L}^{2}\leq 2\lambda_{n}\|\mathbf{x}-\mathbf{x}_{c}\|^{2}+2\lambda_{n}\|\mathbf{x}_{\perp}\|^{2}.$
Define the Lypanuov function of (21) $V:=\chi_{0}V_{1}+V_{2}$ with $\chi_{0}=\frac{4\lambda_{n}c_{4}\sqrt{n}}{\lambda_{2}}$. If we let $\alpha\leq\mathrm{min}\{\frac{2c_{3}}{9\lambda_{n}c_{4}\sqrt{n}},\frac{2c_{3}}{3\lambda_{n}}\}$, then with (22) and (24), we have
$\dot{V}\leq-\frac{\alpha}{4}\chi_{0}\lambda_{2}\|\mathbf{x}_{\perp}\|^{2}-\frac{c_{3}}{3}\|\mathbf{x}-\mathbf{x}_{c}\|^{2}.$ (25)
With (23), we have $V\geq\frac{\chi_{0}}{2}\|\mathbf{x}_{\perp}\|^{2}+c_{1}\|\mathbf{x}-\mathbf{x}_{c}\|^{2},$ then $\|\mathbf{x}_{\perp}(t)\|^{2}=\mathcal{O}(e^{-\gamma t})$, where $\gamma=\mathrm{min}\{\frac{\alpha\lambda_{2}}{2},\frac{c_{3}}{3c_{1}}\}$. The theorem is proved.
.4 Proof of Theorem 3
Flow (10) can be written in a tight form as
$\dot{\mathbf{x}}$ $=-\mathcal{L}{\mathbf{x}_{c}}-\beta\mathbf{v}-\eta\mathbf{F}(\mathbf{x}),$ (26)
$\dot{\mathbf{v}}$ $={\beta\mathcal{L}\mathbf{x}_{c}},$
$\mathbf{x}_{c}$ $=\mathcal{C}(\mathbf{x},t),$
where $\mathbf{v}:=[\mathbf{v}_{1}^{T},\dots,\mathbf{v}_{n}^{T}]^{T}$ and $\mathbf{F}(\mathbf{x}):=[\nabla f_{1}^{T}(\mathbf{x}_{1}),\dots,\nabla f_{n}^{T}(\mathbf{x}_{n})]^{T}$.
As $f(x)$ is strongly convex, there exists a unique $s^{\ast}\in\mathbb{R}^{d}$ that $\nabla f(s^{\ast})=\mathbf{0}_{d}$, i.e., $\mathbb{I}^{T}\mathbf{F}(\mathbf{s})=\mathbf{0}_{d}$, where $\mathbf{s}:=\sqrt{n}\mathbb{I}s^{\ast}$. It can be noticed that $\mathbf{x}^{\ast}=\mathbf{s}$, and $\mathbf{v}^{\ast}=-\frac{\eta\mathbf{F}(\mathbf{x}^{\ast})}{\beta}$ is the equilibrium point of the system (26). We introduce the state error by defining $\overline{\mathbf{x}}:=\mathbf{x}-\mathbf{s}$, $\overline{\mathbf{v}}:=\mathbf{v}-\mathbf{v}^{\ast}$ Taking the time derivative of the state errors along (26) yields
$\dot{\overline{\mathbf{x}}}$ $=-\mathcal{L}\mathbf{x}_{c}-\beta\overline{\mathbf{v}}-\eta\tilde{\mathbf{F}}(\overline{\mathbf{x}}),$ (27)
$\dot{\overline{\mathbf{v}}}$ $=\beta\mathcal{L}\mathbf{x}_{c},$
$\mathbf{x}_{c}$ $=\mathcal{C}(\mathbf{x},t),$
where $\tilde{\mathbf{F}}(\overline{\mathbf{x}}):=\mathbf{F}(\overline{\mathbf{x}}+\mathbf{s})-\mathbf{F}(\mathbf{s})$.
We decompose $\overline{\mathbf{x}}$ and $\overline{\mathbf{v}}$ by defining $\mathbf{\overline{x}}_{\perp}:=\mathbb{S}^{T}\overline{\mathbf{x}}$, $\mathbf{\overline{x}}_{\parallel}:=\mathbb{I}^{T}\overline{\mathbf{x}}$, $\mathbf{\overline{v}}_{\perp}:=\mathbb{S}^{T}\overline{\mathbf{v}}$ and $\mathbf{\overline{v}}_{\parallel}:=\mathbb{I}^{T}\overline{\mathbf{v}}$.
It can be noticed that the convergence of $\overline{\mathbf{x}}(t)$ and $\overline{\mathbf{v}}(t)$ can be shown as if $\mathbf{\overline{x}}_{\parallel}(t)$, $\mathbf{\overline{x}}_{\perp}(t)$, $\mathbf{\overline{v}}_{\perp}(t)$ and $\mathbf{\overline{v}}_{\perp}(t)$ converge to the zero equilibrium by (17).
By the initial condition, we know that $\mathbb{I}^{T}\mathbf{v}(0)=\mathbf{0}_{d}$. With systems (26) and (16), we conclude that $\mathbb{I}^{T}\mathbf{v}(t)=\mathbf{0}_{d}$, then
$\mathbf{\overline{v}}_{\parallel}(t)=\mathbb{I}^{T}\overline{\mathbf{v}}(t)=\mathbb{I}^{T}(\mathbf{v}(t)-\frac{\eta\mathbf{F}(\mathbf{s})}{\beta})=\mathbf{0}_{d}.$ (28)
By (16) and (28), the system (27) becomes
$\mathbf{\dot{\overline{x}}}_{\perp}$ $=-\mathbb{S}^{T}\mathcal{L}\mathbf{x}_{c}-\beta\mathbf{\overline{v}}_{\perp}-\eta\mathbb{S}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}}),$ (29)
$\mathbf{\dot{\overline{x}}}_{\parallel}$ $=-\eta\mathbb{I}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}}),$
$\mathbf{\dot{\tilde{v}}}_{\perp}$ $=\beta\mathbb{S}^{T}\mathcal{L}\mathbf{x}_{c},$
$\mathbf{x}_{c}$ $=\mathcal{C}(\mathbf{x},t)-\mathcal{C}(\mathbb{I}\mathbf{\overline{x}}_{\parallel}+\mathbf{s},t),$
where the fact $\mathcal{L}\mathcal{C}(\mathbb{I}\overline{\mathbf{x}}+\mathbf{s},t)=\mathbf{0}_{nd}$ is used.
Let $\mathbf{z}:=\frac{1}{\beta}\mathbf{\overline{v}}_{\perp}+\mathbf{\overline{x}}_{\perp}$. The system (29) then becomes
$\mathbf{\dot{\overline{x}}}_{\perp}$ $=-\mathbb{S}^{T}\mathcal{L}\mathbf{x}_{c}+\beta^{2}\mathbf{\overline{x}}_{\perp}-\beta^{2}\mathbf{z}-\eta\mathbb{S}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}}),$ (30)
$\mathbf{\dot{\overline{x}}}_{\parallel}$ $=-\eta\mathbb{I}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}}),$
$\dot{\mathbf{z}}$ $=-\beta^{2}\mathbf{z}+\beta^{2}\mathbf{\overline{x}}_{\perp}-\eta\mathbb{S}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}}),$
$\mathbf{x}_{c}$ $=\mathcal{C}(\mathbf{x},t)-\mathcal{C}(\mathbb{I}\mathbf{\overline{x}}_{\parallel}+\mathbf{s},t).$
[…]
Define $V_{1}(\mathbf{\overline{x}}_{\perp},\mathbf{z})=\frac{1}{2}(\|\mathbf{\overline{x}}_{\perp}\|^{2}+\|\mathbf{z}\|^{2})$, then we have
$\dot{V}_{1}$ $\leq-\mathbf{\overline{x}}_{\perp}^{T}\mathbb{S}\mathcal{L}\mathbf{x}_{c}-\beta^{2}\|\mathbf{z}\|^{2}+\beta^{2}\|\mathbf{\overline{x}}_{\perp}\|^{2}$ (31)
$\quad-\eta\mathbf{z}^{T}\mathbb{S}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}})-\eta\mathbf{\overline{x}}_{\perp}^{T}\mathbb{S}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}})]$
$\leq L_{c}\lambda_{n}\|\mathbf{\overline{x}}_{\perp}\|^{2}-(\beta^{2}-\frac{\eta}{2})\|\mathbf{z}\|^{2}$
$\quad+(\beta^{2}+\frac{\eta}{2}+{\eta}L_{f}^{2})\|\mathbf{\overline{x}}_{\perp}\|^{2}+\eta L_{f}^{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2},$
where the second inequality is obtained by the fact
$\|\mathbf{x}_{c}\|$ $=\|\mathcal{C}(\mathbf{x},t)-\mathcal{C}(\mathbb{I}\mathbf{\overline{x}}_{\parallel}+\mathbf{s},t)\|$ (32)
$\leq L_{c}\|\mathbf{x}-\mathbb{I}\mathbf{\overline{x}}_{\parallel}-\mathbf{s}\|\leq L_{c}\|\mathbf{\overline{x}}_{\perp}\|$
from properties of $\mathbf{C}$ and
$\|\tilde{\mathbf{F}}(\overline{\mathbf{x}})\|^{2}\leq L^{2}_{f}(\|\mathbf{\overline{x}}_{\perp}\|^{2}+\|\mathbf{\overline{x}}_{\parallel}\|^{2}).$ (33)
Define $V_{2}(\mathbf{\overline{x}}_{\perp},t):=V_{e}(\mathbf{\overline{x}}_{\perp},t)$, where $V_{e}$ is defined in (7), then we have
$\dot{V}_{2}$ $=\frac{\partial V}{\partial t}-\frac{\partial V}{\partial\mathbf{\overline{x}}_{\perp}}\Lambda\mathcal{C}(\mathbf{\overline{x}}_{\perp},t)+\frac{\partial V}{\partial\mathbf{\overline{x}}_{\perp}}[\Lambda\mathcal{C}(\mathbf{\overline{x}}_{\perp},t)$ (34)
$\quad-\mathbb{S}^{T}\mathcal{L}\mathbf{\mathcal{C}}(\mathbf{x},t)]+\frac{\partial V_{2}}{\partial\mathbf{x}_{\perp}}(\beta^{2}\mathbf{\overline{x}}_{\perp}-\beta^{2}\mathbf{z}-\eta\mathbb{S}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}}))$
$\leq-(c^{\prime}_{3}-c_{4}\beta^{2}-c_{4}\beta^{2}/r-c_{4}\eta/r-c_{4}\eta L_{f}^{2}r)\|\mathbf{\overline{x}}_{\perp}\|^{2}$
$\quad+c_{4}\beta^{2}r\|\mathbf{z}\|^{2}+c_{4}\eta rL_{f}^{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2},$
where we let $\delta\leq\frac{c_{3}}{c_{4}\lambda_{n}}$ and then $c^{\prime}_{3}:=c_{3}-c_{4}\delta\lambda_{n}>0$, and the last inequality is obtained by (19) with the fact $\mathbb{S}^{T}\mathbf{x}=\mathbb{S}^{T}\overline{\mathbf{x}}=\mathbf{\overline{x}}_{\perp}$, (33) and Young’s Inequality, where $r>0$ is a parameter that will be determined later.
Define $V_{3}(\mathbf{\overline{x}}_{\parallel}):=\frac{1}{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}$, as $f(x)$ is $\mu$-strongly convex, then we have
$\dot{V}_{3}$ $=-\eta\mathbf{\overline{x}}_{\parallel}^{T}\mathbb{I}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}})$ (35)
$=-\eta\mathbf{\overline{x}}_{\parallel}^{T}\mathbb{I}^{T}[\mathbf{F}(\overline{\mathbf{x}}+\mathbf{s})-\mathbf{F}(\mathbb{S}\mathbf{\overline{x}}_{\perp}+\mathbf{s})$
$\quad+\mathbf{F}(\mathbb{S}\mathbf{\overline{x}}_{\perp}+\mathbf{s})-\mathbf{F}(\mathbf{s})]$
$\leq-\frac{\eta\mu_{n}}{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}+\frac{\eta}{2\mu_{n}}L_{f}^{2}\|\mathbf{\overline{x}}_{\perp}\|^{2},$
where the last inequality is obtained by Property i). and ii). of $f(x)$ in Assumption 1 with $\mu_{n}:=\frac{\mu}{n}$.
We introduce some positive parameters that have nothing to do with $\beta$, $r$ and $\eta$.
$\chi_{0}=2L_{c}\lambda_{n}/c^{\prime}_{3},\quad\chi_{1}=\frac{4L_{f}^{2}}{\mu_{n}}+\frac{4\chi_{0}c_{4}L_{f}^{2}}{\mu_{n}},$
$\xi_{1}=\frac{3}{2}+L_{f}^{2}+\chi_{0}(c_{4}+c_{4}L_{f}^{2})+\chi_{1}\frac{L_{f}^{2}}{2\mu_{n}},$
$\xi_{2}=2\chi_{0}c_{4},\quad\xi_{3}=\chi_{0}c_{4},\quad\xi_{4}=\frac{1}{2}.$
We define the Lyapunov functions of the system (30) $V:=V_{1}+\chi_{0}V_{2}+\chi_{1}V_{3}$. It is easy to prove that $V$ is positive definite. In fact,
$V\geq(\frac{1}{2}+\chi_{0}c_{1})\|\mathbf{\overline{x}}_{\perp}\|^{2}+\frac{1}{2}\|\mathbf{z}\|^{2}+\frac{\chi_{1}}{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}.$ (36)
By (31), (34), (35), we know $\dot{V}$ is negative when we choose $r=\mathrm{min}\{\frac{1}{4\xi_{3}},1\}$, $\beta^{2}\leq\mathrm{min}\{\frac{\chi_{0}c^{\prime}_{3}}{8\xi_{1}},\frac{\chi_{0}c^{\prime}_{3}r}{8\xi_{2}}\}$, $\eta\leq\mathrm{min}\{\beta^{2},\frac{\beta^{2}}{4\xi_{4}}\}$. With (36), we have
$\dot{V}\leq-\gamma V,\quad\gamma=\mathrm{min}\{\frac{\chi_{0}c^{\prime}_{3}}{2(1+2\chi_{0}c_{1})},\beta^{2},\eta\frac{\mu_{n}}{2}\}.$
With the definition of $V$, we derive $\|\overline{\mathbf{x}}(t)\|^{2}=\mathcal{O}(e^{-\gamma t})$. With the definition $\overline{\mathbf{x}}=\mathbf{x}-\mathbf{s}$ before, we know that $\mathbf{x}_{i}(t)$ in Flow (10) converge exponentially to the optimal solution $s^{\ast}$ with the SST compressor.
.5 Proof of Theorem 4
As analyzed in Appendix .3, Flow (11) satisfies that for every $i\in\mathrm{V}$, ${\hat{\mathbf{x}}}^{j}_{i}(0)={\hat{\mathbf{x}}}^{j^{\prime}}_{i}(0),\forall j,j^{\prime}\in\mathrm{V}$. Then Flow (11) can be written as
$\dot{\mathbf{x}}$ $=-\alpha\mathcal{L}{\mathbf{x}_{c}}-\beta\mathbf{v}-\eta\mathbf{F}(\mathbf{x}),$ (37)
$\dot{\mathbf{v}}$ $={\beta\mathcal{L}\mathbf{x}_{c}},$
$\dot{\mathbf{x}}_{c}$ $=\mathcal{C}(\mathbf{x}-\mathbf{x}_{c},t).$
We carry out a similar analysis process as in Appendix .4 with $\mathbf{z}:=\frac{\alpha}{\beta}\mathbf{\overline{v}}_{\perp}+\mathbf{\overline{x}}_{\perp}$, and the system (37) becomes
$\mathbf{\dot{\tilde{x}}}_{\perp}$ $=-\alpha\mathbb{S}^{T}\mathcal{L}\overline{\mathbf{x}}_{c}+\beta^{2}_{\alpha}\mathbf{\overline{x}}_{\perp}-\beta^{2}_{\alpha}\mathbf{z}-\eta\mathbb{S}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}}),$ (38)
$\mathbf{\dot{\tilde{x}}}_{\parallel}$ $=-\eta\mathbb{I}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}}),$
$\dot{\mathbf{z}}$ $=-\beta^{2}_{\alpha}\mathbf{z}+\beta^{2}_{\alpha}\mathbf{\overline{x}}_{\perp}-\eta\mathbb{S}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}}),$
$\dot{\tilde{\mathbf{x}}}_{c}$ $=\mathcal{C}(\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c},t),$
where $\overline{\mathbf{x}}_{c}:=\mathbf{x}_{c}-\mathbf{s}$ and $\beta^{2}_{\alpha}:=\beta^{2}/\alpha$.
Define $V_{1}(\mathbf{\overline{x}}_{\perp},\mathbf{z})=\frac{1}{2}(\|\mathbf{\overline{x}}_{\perp}\|^{2}+\|\mathbf{z}\|^{2})$, then we have
$\dot{V}_{1}$ $\leq-\alpha\mathbf{\overline{x}}_{\perp}^{T}\mathbb{S}\mathcal{L}\overline{\mathbf{x}}_{c}-\beta^{2}_{\alpha}\|\mathbf{z}\|^{2}+\beta^{2}_{\alpha}\|\mathbf{\overline{x}}_{\perp}\|^{2}$ (39)
$\quad-\eta\mathbf{z}^{T}\mathbb{S}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}})-\eta\mathbf{\overline{x}}_{\perp}^{T}\mathbb{S}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}})]$
$\leq-\frac{\alpha\lambda_{2}}{2}\|\mathbf{\overline{x}}_{\perp}\|^{2}-(\beta^{2}_{\alpha}-\frac{\eta}{2})\|\mathbf{z}\|^{2}$
$\quad+(\beta^{2}_{\alpha}+\frac{\eta}{2}+{\eta}L_{f}^{2})\|\mathbf{\overline{x}}_{\perp}\|^{2}+\eta L_{f}^{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}$
$\quad+\frac{\alpha\lambda_{n}}{2}\|\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c}\|^{2},$
where the second inequality is obtained by the fact
$-\mathbf{\overline{x}}_{\perp}^{T}\mathbb{S}^{T}\mathcal{L}\overline{\mathbf{x}}_{c}$ $\leq-\frac{1}{2}\lambda_{2}(\|\mathbf{\overline{x}}_{\perp}\|^{2})+\frac{1}{2}\lambda_{n}(\|\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c}\|^{2}).$ (40)
For $V_{2}(\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c},t)$, which is defined in Appendix .3, we have
$\dot{V}_{2}$ $\leq-c_{3}\|\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c}\|^{2}$ (41)
$\quad+c_{4}\sqrt{n}\|\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c}\|\|\alpha\mathcal{L}\overline{\mathbf{x}}_{c}+\beta^{2}_{\alpha}\mathbb{S}\mathbf{z}_{k}-\beta_{\alpha}^{2}\mathbb{S}\mathbf{\overline{x}}_{\perp}+\eta\tilde{\mathbf{F}}(\overline{\mathbf{x}})\|$
$\leq-[c_{3}-c_{4}\sqrt{n}\big{(}\frac{\alpha}{r}+\frac{2\beta_{\alpha}^{2}}{r}+\frac{\eta}{r}+2\alpha r\lambda_{n}^{2}\big{)}]\|\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c}\|^{2}$
$\quad+c_{4}\sqrt{n}\beta^{2}_{\alpha}r\|\mathbf{z}\|^{2}+(2c_{4}\sqrt{n}\alpha r\lambda_{n}^{2}+c_{4}\beta^{2}_{\alpha}r)\|\mathbf{\overline{x}}_{\perp}\|^{2}$
$\quad+c_{4}\sqrt{n}\eta rL_{f}^{2}\|\mathbf{\overline{x}}_{\perp}\|^{2}+c_{4}\sqrt{n}\eta rL_{f}^{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2},$
where the first inequality is obtained by $\overline{\mathbf{v}}=\mathbb{S}\mathbf{\overline{v}}_{\perp}$ and (23), and the last inequality is obtained by (33), the fact
$\|\mathbf{x}_{c}\|_{L^{2}}^{2}$ $\leq 2\lambda_{n}^{2}\|\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c}\|^{2}+2\lambda_{n}^{2}\|\mathbf{\overline{x}}_{\perp}\|^{2},$ (42)
and Young’s Inequality, where $r>0$ is a parameter that will be determined later.
Define $V_{3}(\mathbf{\overline{x}}_{\parallel}):=\frac{1}{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}$, then (35) still holds.
We introduce some positive parameters that have nothing to do with $\alpha$, $\beta$, $r$ and $\eta$.
$\quad\chi_{1}=\frac{4L_{f}^{2}}{\mu_{n}}+\frac{4c_{4}\sqrt{n}L_{f}^{2}}{\mu_{n}},\quad\xi_{1}=\frac{\lambda_{2}}{2},$
$\quad\xi_{2}=\frac{3}{2}+L_{f}^{2}+c_{4}\sqrt{n}+c_{4}\sqrt{n}L_{f}^{2}+\chi_{1}\frac{L_{f}^{2}}{2\mu_{n}},$
$\quad\xi_{3}=2c_{4}\sqrt{n}\lambda_{n}^{2},\quad\xi_{4}=\frac{1}{2},\quad\xi_{5}=c_{4}\sqrt{n},$
$\quad\xi_{6}=4c_{4}\sqrt{n},\quad\xi_{7}=\frac{\lambda_{n}}{2}+2c_{4}\sqrt{n}\lambda_{n}^{2}.$
We define the Lyapunov functions of the system (38) $V:=V_{1}+V_{2}+\chi_{1}V_{3}$. It is easy to prove that $V$ is positive definite. In fact,
$V\geq\frac{1}{2}\|\mathbf{\overline{x}}_{\perp}\|^{2}+\frac{1}{2}\|\mathbf{z}\|^{2}+\frac{\chi_{1}}{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}+c_{1}\|\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c}\|.$ (43)
By (39), (41) and (35), we know $\dot{V}$ is negative when we choose $r=\mathrm{min}\{\frac{\xi_{1}}{4\xi_{3}},\frac{1}{4\xi_{5}},1\}$, $\alpha\leq\mathrm{min}\{\frac{c_{3}r}{4\xi_{6}},\frac{c_{3}}{4\xi_{7}}\}$, $\beta^{2}\leq\mathrm{min}\{{\alpha^{2}},\frac{\xi_{1}\alpha^{2}}{4\xi_{2}}\}$, $\eta\leq\mathrm{min}\{\beta_{\alpha}^{2},\frac{1}{4\xi_{4}}\}$. With (43), we have
$\dot{V}\leq-\gamma V,\quad\gamma=\mathrm{min}\{\xi_{1}\alpha,\beta^{2}_{\alpha},\eta\frac{\mu_{n}}{2},\frac{c_{3}}{2c_{1}}\}.$
With the definition of $V$, we derive $\|\overline{\mathbf{x}}(t)\|^{2}=\mathcal{O}(e^{-\gamma t})$. With the definition $\overline{\mathbf{x}}=\mathbf{x}-\mathbf{s}$ before, we know that $\mathbf{x}_{i}(t)$ in Flow (11) converge exponentially to the optimal solution $s^{\ast}$ with the ST compressor.
.6 Proof of Proposition 2
Proof of a). The proof is similar to that of Proposition 1 as ${\mathbf{x}}_{e}(t+1)=\mathbf{x}_{e}(t)-\kappa_{0}\bm{\psi}(t)\bm{\psi}(t)^{T}\mathbf{x}_{e}(t)$ is globally linearly stable at the zero equilibrium for any $\kappa_{0}\leq\kappa_{0}^{\ast}$ with some $\kappa_{0}^{\ast}>0$ under the condition in discrete-time cases by recalling [43].
Proof of b). Next, we prove that $\mathbf{C}_{2}$ is the ST compressor in discrete time by proving the system $\mathbf{x}_{e}(t+1)=\mathbf{x}_{e}(t)-\kappa_{0}\mathbf{C}_{2}(\mathbf{x}_{e}(t))$ is linearly stable at the zero equilibrium with $\kappa_{0}=\frac{1}{p}$. By (13),
$\|\mathbf{x}_{e}(t+1)\|^{2}-\|\mathbf{x}_{e}(t)\|^{2}$ $=-\frac{2\mathbf{C}_{2}(\mathbf{x}_{e}(t))^{T}\mathbf{x}_{e}(t)}{p}+\|\frac{\mathbf{C}_{2}(\mathbf{x}_{e}(t))}{p}\|^{2}$
$\leq-\varphi\|\mathbf{x}_{e}(t)\|^{2}.$
Thus $\mathbf{x}_{e}$-system is linearly stable at the zero equilibrium with $\varphi\in(0,1]$. With (14), the proof is complete.
.7 Proof for Theorem 5
Similar to the proof of continuous-time form in Appendix .4, DPD-DC can be written in a tight form. We introduce the state error by defining $\overline{\mathbf{x}}:=\mathbf{x}-\mathbf{s}$, $\overline{\mathbf{v}}:=\mathbf{v}-\mathbf{v}^{\ast}$. Then decompose $\overline{\mathbf{x}}$ and $\overline{\mathbf{v}}$ by defining $\mathbf{\overline{x}}_{\perp}:=\mathbb{S}^{T}\overline{\mathbf{x}}$, $\mathbf{\overline{x}}_{\parallel}:=\mathbb{I}^{T}\overline{\mathbf{x}}$, $\mathbf{\overline{v}}_{\perp}:=\mathbb{S}^{T}\overline{\mathbf{v}}$ and $\mathbf{\overline{v}}_{\parallel}:=\mathbb{I}^{T}\overline{\mathbf{v}}$. The convergence of $\overline{\mathbf{x}}(t)$ and $\overline{\mathbf{v}}(t)$ can be shown as if $\mathbf{\overline{x}}_{\parallel}(t)$, $\mathbf{\overline{x}}_{\perp}(t)$, $\mathbf{\overline{v}}_{\perp}(t)$ and $\mathbf{\overline{v}}_{\perp}(t)$ converge to the zero equilibrium. Also, we can conclude (28).
Let $\mathbf{z}:=\frac{1}{\beta}\mathbf{\overline{v}}_{\perp}+\mathbf{\overline{x}}_{\perp}$. With (30), the system becomes
[…]
It can be noticed that the linear stability of the system DPD-DC and the system (44) are equal.
Define $V_{1,t}(\mathbf{\overline{x}}_{\perp},\mathbf{z})=\frac{1}{2}(\|\mathbf{\overline{x}}_{\perp}\|^{2}+\|\mathbf{z}\|^{2})$, then
$\Delta{V}_{1,t}$ $\leq(L_{c}\lambda_{n}\kappa_{0}+2L_{c}^{2}\lambda^{2}_{n}\kappa_{0}^{2})\|\mathbf{\overline{x}}_{\perp}\|^{2}$ (45)
$\quad+\kappa[-(\beta^{2}-\frac{\eta}{2})\|\mathbf{z}\|^{2}+(\beta^{2}+\frac{\eta}{2}+{\eta}L_{f}^{2})\|\mathbf{\overline{x}}_{\perp}\|^{2}$
$\quad+\eta L_{f}^{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}]+\frac{1}{2}\kappa^{2}[(7\eta^{2}L_{f}^{2}+7\beta^{4})\|\mathbf{\overline{x}}_{\perp}\|^{2}$
$\quad+7\beta^{4}\|\mathbf{z}\|^{2}+7\eta^{2}L_{f}^{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}],$
where the inequality is obtained by (32) and (33).
Before we introduce the second Lyapunov function, we will show that the following system,
${\mathbf{y}}_{e}(t+1)=\mathbf{y}_{e}(t)-\kappa_{0}\mathbb{S}^{T}\mathcal{L}\mathbf{\mathcal{C}}(\mathbf{x}_{e}(t),t),$ (46)
where $\mathbf{y}_{e}\in\mathbb{R}^{(n-1)d}$, $\mathbf{x}_{e}\in\mathbb{R}^{nd}$ and ${\mathbf{y}}_{e}=\mathbb{S}^{T}{\mathbf{x}}_{e}$, achieves linear convergence at the zero equilibrium for some positive $\kappa_{0},\delta$ in Theorem 5.
By definition of $\mathbf{C}(\mathbf{x}_{e},t)$, it is easy to find the following system achieves linear convergence at the zero equilibrium if $\kappa_{0}\leq\kappa_{0}^{\ast}/\lambda_{n}$,
${\mathbf{y}}_{e}(t+1)=\mathbf{y}_{e}(t)-\kappa_{0}\Lambda\mathcal{C}(\mathbf{y}_{e}(t),t).$
Then there exist positive constants $C$, $\gamma_{D}<1$. For any $t$ and $N\in\mathbb{N}_{+}$, the solution satisfies
$(\|\mathbf{y}_{e}(t+N)\|^{2})\leq C(\|\mathbf{y}_{e}(t)\|^{2})\gamma_{D}^{N}.$
We assume $\phi_{t}^{t+T}(\mathbf{y}_{e}(t))$ is the state of the system $\mathbf{y}_{e}(t+1)=\mathbf{y}_{e}(t)-\kappa_{0}\Lambda\mathcal{C}(\mathbf{y}_{e}(t),t)$ in $t+N$ moment with the state in $t$ moment is $\mathbf{y}_{e}(t)$. It is easy to verify that
$\|\phi_{t}^{t+N}(\mathbf{y})\|^{2}$ $\leq L_{\phi}\|\mathbf{y}\|^{2}$
for any $\mathbf{y}\in\mathbb{R}^{(n-1)d}$ and some $L_{\phi}>0$ by the property of compressor $\mathbf{C}$.
We find the Lyapunov function $V_{e}(\mathbf{y}_{e},t)=\sum_{j=0}^{N-1}\|\phi_{t}^{t+j}(\mathbf{y}_{e})\|^{2}$ satisfies
$\quad c_{1}\|\mathbf{y}_{e}\|^{2}\leq V_{e,t}\leq c_{2}\|\mathbf{y}_{e}\|^{2}$ (47)
for $c_{1}=1,c_{2}=NL_{\phi}$.
Moreover, we have
$\Delta V_{e,t}$ $=\sum_{j=1}^{N}\|\phi_{t+1}^{t+j}(\mathbf{y}_{e}(t+1))\|^{2}-\sum_{j=0}^{N-1}\|\phi_{t}^{t+j}(\mathbf{y}_{e}(t))\|^{2}$
$=\|\mathbf{y}_{e}(t+N)\|^{2}-\|\mathbf{y}_{e}(t)\|^{2}$
$\leq-(1-C\gamma_{D}^{N})\|\mathbf{y}_{e}(t)\|^{2}\leq-c_{3}\|\mathbf{y}_{e}(t)\|^{2}.$
We choose a $N\in\mathbb{N}_{+}$ large enough and then $c_{3}=1-C\gamma_{D}^{N}>0$, i.e.,
$\quad\sum_{j=1}^{N}\|\phi_{t+1}^{t+j}(\mathbf{y}_{e}-\kappa_{0}\Lambda\mathcal{C}(\mathbf{y}_{e},t))\|^{2}-\sum_{j=0}^{N-1}\|\phi_{t}^{t+j}(\mathbf{y}_{e})\|^{2}$ (48)
$\leq-c_{3}\|\mathbf{y}_{e}\|^{2}.$
Besides,
$\|\mathbf{y}_{e}-\kappa_{0}\Lambda\mathcal{C}(\mathbf{y}_{e},t)\|^{2}\leq\theta\|\mathbf{y}_{e}\|^{2},$ (49)
for $\theta=2+2L_{c}^{2}\kappa_{0}^{2}\lambda_{n}^{2}>0$ by the property of $\mathbf{C}$.
For the system (46), we apply the Lyapunov function $V_{e}(\mathbf{y}_{e},t)$ and obtain the difference of $V_{e}(\mathbf{y}_{e},t)$
$\quad\sum_{j=1}^{N}\|\phi_{t+1}^{t+j}(\mathbf{y}_{e}-\kappa_{0}\mathbb{S}^{T}\mathcal{L}\mathbf{\mathcal{C}}(\mathbf{x}_{e}(t),t)\|^{2}-\sum_{j=0}^{N-1}\|\phi_{t}^{t+j}(\mathbf{y}_{e})\|^{2}$ (50)
$\leq-c_{3}\|\mathbf{y}_{e}\|^{2}+c_{4}\kappa_{0}^{2}\lambda_{n}^{2}\|\mathcal{C}(\mathbb{S}^{T}\mathbf{x}_{e}(t),t)-\mathbb{S}^{T}\mathcal{C}(\mathbf{x}_{e}(t),t)\|^{2}$
$\quad+2c_{4}\kappa_{0}\lambda_{n}\|\mathbf{y}_{e}\|\|\mathcal{C}(\mathbb{S}^{T}\mathbf{x}_{e}(t),t)-\mathbb{S}^{T}\mathcal{C}(\mathbf{x}_{e}(t),t)\|$
$\leq-(c_{3}-2c_{4}\kappa_{0}\lambda_{n}\delta-c_{4}\kappa_{0}^{2}\lambda_{n}^{2}\delta^{2})\|\mathbf{y}_{e}\|^{2}$
for $c_{4}:=NL_{\phi}\theta$, where the first inequality is obtained by (LABEL:eq:DPD.a.Vec3) and the second inequality is obtained by (8). It is obvious that for $0<\delta<\frac{c_{4}+\sqrt{c_{4}^{2}+c_{3}c_{4}}}{c_{4}\kappa_{0}\lambda_{n}}$, the difference of $V_{e}(\mathbf{y}_{e},t)$ is negative definite with $c_{3}^{\prime}:=c_{3}-2c_{4}\kappa_{0}\lambda_{n}\delta-c_{4}\kappa_{0}^{2}\lambda_{n}^{2}\delta^{2}>0$. Thus the system (46) achieves linear convergence at the zero equilibrium.
Next we continue to choose the Lyapunov function by defining $V_{2,t}:=V_{e}(\mathbf{\overline{x}}_{\perp},t)$, then
$\Delta V_{2,t}$ $=\sum_{j=1}^{N}\|\phi_{t+1}^{t+j}(\mathbf{\overline{x}}_{\perp}(t+1))\|^{2}-\sum_{j=0}^{N-1}\|\phi_{t}^{t+j}(\mathbf{\overline{x}}_{\perp}(t))\|^{2}$ (51)
$\leq-(c^{\prime}_{3}-\kappa c_{4}\beta^{2}-\kappa c_{4}\beta^{2}/r-\kappa c_{4}\eta/r-\kappa c_{4}\eta rL_{f}^{2})\|\mathbf{\overline{x}}_{\perp}\|^{2}$
$\quad+\kappa[c_{4}\beta^{2}r\|\mathbf{z}\|^{2}+c_{4}\eta rL_{f}^{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}]$
$\quad+\kappa^{2}NL_{\phi}[(3\beta^{4}+3\eta^{2}L_{f}^{2})\|\mathbf{\overline{x}}_{\perp}\|^{2}$
$\quad+3\beta^{4}\|\mathbf{z}\|^{2}+3\eta^{2}L_{f}^{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}],$
where the inequality is obtained by (49), (LABEL:eq:DPD.a.Ve'), (33) and Young’s Inequality, where $r>0$ is a parameter that will be determined later.
Define $V_{3,t}(\mathbf{\overline{x}}_{\parallel}):=\frac{1}{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}$. With (35), we have
$\Delta V_{3,t}$ $=\frac{1}{2}(\mathbf{\overline{x}}_{\parallel}-\kappa\eta\mathbb{I}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}}))^{T}(\mathbf{\overline{x}}_{\parallel}-\kappa\eta\mathbb{I}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}}))-\frac{1}{2}\mathbf{\overline{x}}_{\parallel}^{T}\mathbf{\overline{x}}_{\parallel}$ (52)
$\leq\kappa\big{(}-\eta\frac{\mu_{n}}{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}+\eta\frac{1}{2\mu_{n}}L_{f}^{2}\|\mathbf{\overline{x}}_{\perp}\|^{2}\big{)}$
$\quad+\frac{1}{2}\kappa^{2}\eta^{2}L_{f}^{2}(\|\mathbf{\overline{x}}_{\perp}\|^{2}+\|\mathbf{\overline{x}}_{\parallel}\|^{2}),$
where the equality is obtained by (33) and (35).
We introduce some parameters $\chi_{0}$, $\chi_{1}$, $\xi_{1}$, $\xi_{2},\dots>0$ that have nothing to do with $\beta$, $r$ and $\eta$, and some parameters $\zeta_{1}$, $\zeta_{2},\dots>0$.
$\quad\chi_{0}=(2L_{c}\lambda_{n}\kappa_{0}+4L_{c}^{2}\lambda_{n}^{2}\kappa_{0}^{2})/c^{\prime}_{3},\quad\chi_{1}=\frac{4L_{f}^{2}}{\mu_{n}}+\frac{4\chi_{0}c_{4}L_{f}^{2}}{\mu_{n}},$
$\quad\xi_{1}=\frac{3}{2}+L_{f}^{2}+\chi_{0}(c_{4}+c_{4}L_{f}^{2})+\chi_{1}\frac{L_{f}^{2}}{2\mu_{n}},$
$\quad\xi_{2}=2\chi_{0}c_{4},\quad\xi_{3}=\chi_{0}c_{4},\quad\xi_{4}=\frac{1}{2},$
$\quad\zeta_{1}=\frac{7}{2}\eta^{2}L_{f}^{2}+\frac{7}{2}\beta^{4}+\chi_{0}NL_{\phi}(3\beta^{4}+3\eta^{2}L_{f}^{2})+\frac{1}{2}\chi_{1}\eta^{2}L_{f}^{2},$
$\quad\zeta_{2}=\frac{7}{2}\beta^{4}+3\chi_{0}NL_{\phi}\beta^{4},$
$\quad\zeta_{3}=\frac{7}{2}\eta^{2}L_{f}^{2}+3\eta^{2}\chi_{0}NL_{\phi}L_{f}^{2}+\frac{1}{2}\eta^{2}\chi_{1}L_{f}^{2}.$
We define the Lyapunov functions of the system (44) $V_{t}:=V_{1,t}+\chi_{0}V_{2,t}+\chi_{1}V_{3,t}$. It is easy to prove that $V$ is positive definite. In fact,
$V_{t}\geq(\frac{1}{2}+\chi_{0}c_{1})\|\mathbf{\overline{x}}_{\perp}\|^{2}+\frac{1}{2}\|\mathbf{z}\|^{2}+\frac{\chi_{1}}{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}.$ (53)
By (45), (51), (52), we know $\Delta V_{t}$ is negative definite when we choose $r=\mathrm{min}\{\frac{1}{4\xi_{3}},1\}$, $\beta^{2}\leq\mathrm{min}\{\frac{\chi_{0}c^{\prime}_{3}}{8\xi_{1}},\frac{\chi_{0}c^{\prime}_{3}r}{8\xi_{2}}\}$, $\eta\leq\mathrm{min}\{\beta^{2},\frac{\beta^{2}}{4\xi_{4}}\}$ and $\kappa\leq\kappa_{1}:=\frac{1}{2}\mathrm{min}\{\frac{\chi_{0}c^{\prime}_{3}}{4\zeta_{1}},\frac{\beta^{2}}{2\zeta_{2}},\eta\frac{\mu_{n}\chi_{1}}{4\zeta_{3}},1\}$. With (53), we have
$\Delta V_{t}\leq-\gamma V_{t},\ \gamma=\frac{1}{2}\kappa\mathrm{min}\{\frac{\chi_{0}c^{\prime}_{3}}{2(1+2\chi_{0}c_{1})},\beta^{2},\eta\frac{\mu_{n}}{2}\}.$
Let $\kappa_{2}:=2/\mathrm{min}\{\frac{\chi_{0}c^{\prime}_{3}}{2(1+2\chi_{0}c_{1})},\beta^{2},\eta\frac{\mu_{n}}{2}\}$. When $\kappa\leq\mathrm{min}\{\kappa_{1},\kappa_{2}\}$, we can derive for some $\gamma\in(0,1)$ , $V_{t}=\mathcal{O}((1-\gamma)^{t})$. With the definition of $V_{t}$, we derive $\|\overline{\mathbf{x}}(t)\|^{2}=\mathcal{O}((1-\gamma)^{t})$. With the definition $\overline{\mathbf{x}}=\mathbf{x}-\mathbf{s}$ before, we know that $\mathbf{x}_{i}(t)$ in DPD-OC converges linearly to the optimal solution $s^{\ast}$ with the SST compressor.
.8 Proof of Theorem 6
Based on the proof of Theorem 5 and Theorem 4, with (37) in mind, the system DPD-OC becomes
$\mathbf{{\overline{x}}}_{\perp}(t+1)$ $=\mathbf{{\overline{x}}}_{\perp}(t)-\kappa\mathbb{S}^{T}\mathcal{L}\overline{\mathbf{x}}_{c}(t)$ (54)
$\quad+\kappa[\beta^{2}\mathbf{\overline{x}}_{\perp}(t)-\beta^{2}\mathbf{z}(t)-\eta\mathbb{S}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}}(t))],$
$\mathbf{{\overline{x}}}_{\parallel}(t+1)$ $=\mathbf{{\overline{x}}}_{\parallel}(t)-\kappa\eta\mathbb{I}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}}(t)),$
${\mathbf{z}(t+1)}$ $=\mathbf{z}(t)+\kappa[-\beta^{2}\mathbf{z}(t)+\beta^{2}\mathbf{\overline{x}}_{\perp}(t)-\eta\mathbb{S}^{T}\tilde{\mathbf{F}}(\overline{\mathbf{x}}(t))],$
$\overline{\mathbf{x}}_{c}(t+1)$ $=\overline{\mathbf{x}}_{c}(t)+\kappa_{0}\mathcal{C}(\overline{\mathbf{x}}(t)-\overline{\mathbf{x}}_{c}(t),t).$
Define $V_{1,t}(\mathbf{\overline{x}}_{\perp},\mathbf{z})=\frac{1}{2}(\|\mathbf{\overline{x}}_{\perp}\|^{2}+\|\mathbf{z}\|^{2})$, then
$\Delta{V}_{1,t}$ $\leq\kappa\frac{1}{2}\lambda_{n}\|\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c}\|^{2}+2\kappa^{2}\lambda_{n}^{2}\|\overline{\mathbf{x}}_{c}\|^{2}$ (55)
$\quad+\kappa[-(\beta^{2}-\frac{\eta}{2})\|\mathbf{z}\|^{2}$
$\quad+(-\frac{1}{2}\lambda_{2}+\beta^{2}+\frac{\eta}{2}+{\eta}L_{f}^{2})\|\mathbf{\overline{x}}_{\perp}\|^{2}+\eta L_{f}^{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}]$
$\quad+\frac{7}{2}\kappa^{2}[(\eta^{2}L_{f}^{2}+\beta^{4})\|\mathbf{\overline{x}}_{\perp}\|^{2}$
$\quad+\beta^{4}\|\mathbf{z}\|^{2}+\eta^{2}L_{f}^{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}]$
where the inequality is obtained by (33) and (40).
Now that $\mathbf{x}_{e}(t+1)-\mathbf{x}_{e}(t)=-\kappa_{0}\mathbf{C}(\mathbf{x}_{e}(t),t)$, where $\mathbf{x}_{e}\in\mathbb{R}^{d}$, achieves linear convergence at the zero equilibrium, then clearly $\mathbf{y}_{e}(t+1)-\mathbf{y}_{e}(t)=-\kappa_{0}\mathcal{C}(\mathbf{y}_{e}(t),t)$, where $\mathbf{y}_{e}\in\mathbb{R}^{nd}$, achieves also. Similar to $V_{e}$ in Appendix .7, with (LABEL:eq:DPD.a.Vec3) in mind, we can prove the Lyapunov function $V_{e}(\mathbf{y}_{e},t)=\sum_{j=0}^{N-1}\|\phi_{t}^{t+j}(\mathbf{y}_{e})\|^{2}$ with some $N>0$ satisfies
$\quad c_{1}\|\mathbf{y}_{e}\|^{2}\leq V_{e,t}\leq c_{2}\|\mathbf{y}_{e}\|^{2}$ (56)
$\quad\sum_{j=1}^{N}\|\phi_{t+1}^{t+j}(\mathbf{y}_{e}-\kappa_{0}\mathcal{C}(\mathbf{y}_{e},t))\|^{2}-\sum_{j=0}^{N-1}\|\phi_{t}^{t+j}(\mathbf{y}_{e})\|^{2}$
$\leq-c_{3}\|\mathbf{y}_{e}\|^{2}$
for $c_{1}=1,c_{2}=NL_{\phi},c_{3}>0$.
Besides,
$\|\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c}+\kappa_{0}\mathcal{C}(\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c},t)\|^{2}\leq\theta\|\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c}\|^{2}$ (57)
for $\theta=2+2L_{c}^{2}\kappa_{0}^{2}>0$ by the property of $\mathbf{C}$. Moreover, (42) still holds.
Define $V_{2}(\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c},t):=V_{e}(\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c},t)$, and we can derive
$\Delta V_{2,t}$ $=\sum_{j=1}^{N}\|\phi_{t+1}^{t+j}(\overline{\mathbf{x}}(t+1)-\overline{\mathbf{x}}_{c}(t+1))\|^{2}$ (58)
$\quad-\sum_{j=0}^{N-1}\|\phi_{t}^{t+j}(\overline{\mathbf{x}}(t)-\overline{\mathbf{x}}_{c}(t))\|^{2})$
$\leq-[c_{3}-\kappa\big{(}c_{4}r+2c_{4}\beta^{2}/r+c_{4}\eta/r$
$\quad+2c_{4}r\lambda_{n}^{2})]\|\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c}\|^{2}+\kappa\big{(}c_{4}\beta^{2}r\|\mathbf{z}\|^{2}$
$\quad+(2c_{4}r\lambda_{n}^{2}+c_{4}\beta^{2}r)\|\mathbf{\overline{x}}_{\perp}\|^{2}+c_{4}\eta rL_{f}^{2}\|\mathbf{\overline{x}}_{\perp}\|^{2}$
$\quad+c_{4}\eta rL_{f}^{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}\big{)}$
$\quad+\kappa^{2}NL_{\phi}\big{(}4\lambda_{n}^{2}\|\overline{\mathbf{x}}_{c}\|^{2}+4\beta^{4}\|\overline{\mathbf{v}}\|^{2}+(4\beta^{4}$
$\quad+4\eta^{2}L_{f}^{2})\|\mathbf{\overline{x}}_{\perp}\|^{2}+4\eta^{2}L_{f}^{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}\big{)}$
for $c_{4}:=NL_{\phi}\theta$, where the inequality is obtained by $\overline{\mathbf{v}}=\mathbb{S}\mathbf{\overline{v}}_{\perp}$, (56), (57), (33), (42) and Young’s Inequality, where $r>0$ is a parameter that will be determined later.
Define $V_{3,t}(\mathbf{\overline{x}}_{\parallel}):=\frac{1}{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}$, then (52) still holds.
We introduce some parameters $\chi_{1}$, $\xi_{1},\dots>0$ that have nothing to do with $\alpha$, $\beta$, $r$ and $\eta$, and some parameters $\zeta_{1}$, $\zeta_{2},\dots>0$.
$\quad\chi_{1}=\frac{4L_{f}^{2}}{\mu_{n}}+\frac{4c_{4}L_{f}^{2}}{\mu_{n}},\quad\xi_{1}=\frac{\lambda_{2}}{2},$
$\quad\xi_{2}=\frac{3}{2}+L_{f}^{2}+c_{4}+c_{4}L_{f}^{2}+\chi_{1}\frac{L_{f}^{2}}{2\mu_{n}},\quad\xi_{3}=2c_{4}\lambda_{n}^{2},$
$\quad\xi_{4}=\frac{1}{2},\quad\xi_{5}=c_{4},\quad\xi_{6}=4c_{4},\quad\xi_{7}=\frac{\lambda_{n}}{2}+2c_{4}\lambda_{n}^{2},$
$\quad\zeta_{1}=\frac{7}{2}\eta^{2}L_{f}^{2}+\frac{7}{2}\beta^{4}+4\lambda_{n}^{2}$
$\qquad\quad+NL_{\phi}(8\lambda_{n}^{2}+4\beta^{4}+4\eta^{2}L_{f}^{2})+\frac{1}{2}\chi_{1}\eta^{2}L_{f}^{2},$
$\quad\zeta_{2}=\frac{7}{2}\beta^{4}+4NL_{\phi}\beta^{4},$
$\quad\zeta_{3}=\frac{7}{2}\eta^{2}L_{f}^{2}+4NL_{\phi}\eta^{2}L_{f}^{2}+\frac{1}{2}\eta^{2}\chi_{1}L_{f}^{2},$
$\quad\zeta_{4}=4\lambda_{n}+8NL_{\phi}\lambda_{n}^{2}.$
We define the Lyapunov functions of the system (38) $V_{t}:=V_{1,t}+V_{2,t}+\chi_{1}V_{3,t}$. It is easy to prove that $V_{t}$ is positive definite. In fact,
$V\geq\frac{1}{2}\|\mathbf{\overline{x}}_{\perp}\|^{2}+\frac{1}{2}\|\mathbf{z}\|^{2}+\frac{\chi_{1}}{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}+c_{1}\|\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c}\|.$ (59)
By (55), (58), (52), we know $\Delta{V_{t}}$ is negative when we choose $r=\mathrm{min}\{\frac{\xi_{1}}{4\xi_{3}},\frac{1}{4\xi_{5}},1\}$, $\kappa\leq\kappa_{1}=\frac{c_{3}}{2\xi_{6}r+2\xi_{7}}$, $\beta^{2}\leq\mathrm{min}\{{1},\frac{\xi_{1}}{4\xi_{2}}\}$, $\eta\leq\mathrm{min}\{\beta^{2},\frac{1}{4\xi_{4}}\}$ and $\kappa\leq\kappa_{2}:=\frac{1}{2}\mathrm{min}\{\frac{\xi_{1}}{2\zeta_{1}},\frac{\beta^{2}}{2\zeta_{2}},\eta\frac{\chi_{1}\mu_{n}}{4\zeta_{3}},\frac{c_{3}}{2\zeta_{4}}\}$. With (59), we have
$\Delta V_{t}\leq-\gamma V_{t},\ \gamma=\frac{1}{2}\kappa\mathrm{min}\{\frac{\lambda_{2}}{2},{\beta^{2}},\eta\frac{\mu_{n}}{2},\frac{c_{3}}{2c_{1}}\}.$
Let $\kappa_{3}:=2/\mathrm{min}\{{\xi_{1}},{\beta^{2}},\eta\frac{\mu_{n}}{2},\frac{c_{3}}{2c_{1}}\}$. When $\kappa\leq\mathrm{min}\{\kappa_{1},\kappa_{2},\kappa_{3}\}$, we can derive for some $\gamma\in(0,1)$ , $V_{t}=\mathcal{O}((1-\gamma)^{t})$. With the definition of $V_{t}$, we derive $\|\overline{\mathbf{x}}(t)\|^{2}=\mathcal{O}((1-\gamma)^{t})$. With the definition $\overline{\mathbf{x}}=\mathbf{x}-\mathbf{s}$ before, we know that $\mathbf{x}_{i}(t)$ in DPD-OC converges linearly to the optimal solution $s^{\ast}$ with the ST compressor.
Appendix A Proof of Theorem 7
The idea of proof is quite similar to that in Appendix .8. We just recalculate $\Delta V_{2,t}$ with stochastic impact while the other proof process is the same.
Now that $\mathbf{x}_{e}(t+1)-\mathbf{x}_{e}(t)=-\kappa_{0}\mathbf{C}(\mathbf{x}_{e}(t),t)$, where $\mathbf{x}_{e}(t)\in\mathbb{R}^{d}$, achieves mean square exponential convergence at the zero equilibrium, then clearly $\mathbf{y}_{e}(t+1)-\mathbf{y}_{e}=-\kappa_{0}\mathcal{C}(\mathbf{y}_{e},t)$, where $\mathbf{y}_{e}\in\mathbb{R}^{nd}$, achieves also. Then there exists positive constants $C$, $\gamma<1$, for any $t$ and $T\in\mathbb{N}_{+}$, the solution satisfies
$\mathbb{E}\|\mathbf{y}_{e}(t+T)\|^{2}\leq C(\|\mathbf{y}_{e}(t)\|^{2})\gamma^{T}.$
Assume $\phi_{t}^{t+N}(\mathbf{y}_{e})$ is the state of system $\mathbf{y}_{e}(t+1)-\mathbf{y}_{e}(t)=-\kappa_{0}\mathcal{C}(\mathbf{y}_{e}(t),t)$ in $t+N$ moment with the state in $t$ moment is $\mathbf{y}_{e}(t)$. It is easy to verified that
${rcl}\mathbb{E}\|\phi_{t}^{t+N}(\mathbf{y})\|^{2}$$\leq$ $L_{\phi}\|\mathbf{y}\|^{2}$
for any $\mathbf{y}\in\mathbb{R}^{(n-1)d}$ and some $L_{\phi}>0$ by property of compressor $\mathbf{C}$.
With (LABEL:eq:DPD.a.Vec3) in mind, we can proof Lyapunov function $V_{e}(\mathbf{y}_{e},t)=\sum_{j=0}^{N-1}\|\phi_{t}^{t+j}(\mathbf{y}_{e})\|^{2}$ with some $N>0$ satisfies
${rcl}$ $c_{1}\|\mathbf{y}_{e}\|^{2}\leq\mathbb{E}(V_{e,t})\leq c_{2}\|\mathbf{y}_{e}\|^{2}$ (60)
$\mathbb{E}\sum_{j=1}^{N}\|\phi_{t+1}^{t+j}(\mathbf{y}_{e}-\kappa_{0}C(\mathbf{y}_{e},t))\|^{2}-\mathbb{E}\sum_{j=0}^{N-1}\|\phi_{t}^{t+j}(\mathbf{y}_{e})\|^{2}$
$\leq$ $-c_{3}\|\mathbf{y}_{e}\|^{2}$
for $c_{1}=1,c_{2}=NL_{\phi},c_{3}>0$.
Besides,
${rcl}\mathbb{E}\|\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c}+\kappa_{0}\mathcal{C}(\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c},t)\|^{2}\leq\theta\|\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c}\|^{2},$ (61)
for $\theta=2+2L^{2}_{c}\kappa_{0}^{2}>0$ by property of $\mathbf{C}$.
Define $V_{2}(\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c},t):=V_{e}(\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c},t)$, with (58) in mind, we can derive
${rcl}\mathbb{E}\Delta V_{2,t}$ $=$ $\mathbb{E}\sum_{j=1}^{N}\|\phi_{t+1}^{t+j}(\overline{\mathbf{x}}(t+1)-\overline{\mathbf{x}}_{c}(t+1))\|^{2}$ (62)
$-\mathbb{E}\sum_{j=0}^{N-1}\|\phi_{t}^{t+j}(\overline{\mathbf{x}}(t)-\overline{\mathbf{x}}_{c}(t))\|^{2})$
$\leq$ $-[c_{3}-\kappa\big{(}c_{4}/r+2c_{4}\beta^{2}/r+c_{4}\eta/r)\|\overline{\mathbf{x}}-\overline{\mathbf{x}}_{c}\|^{2}]$
$+\kappa\big{(}c_{4}r\lambda_{n}^{2}\|\overline{\mathbf{x}}_{c}\|^{2}+c_{4}\beta^{2}r\|\mathbf{z}\|^{2}+c_{4}\beta^{2}r\|\mathbf{\overline{x}}_{\perp}\|^{2}+$
$c_{4}\eta rL_{f}^{2}\|\mathbf{\overline{x}}_{\perp}\|^{2}+c_{4}\eta rL_{f}^{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}\big{)}$
$+\kappa^{2}NL_{\phi}\big{(}4\lambda_{n}^{2}\|\overline{\mathbf{x}}_{c}\|^{2}+4\beta^{4}\|\overline{\mathbf{v}}\|^{2}+(4\beta^{4}$
$+4\eta^{2}L_{f}^{2})\|\mathbf{\overline{x}}_{\perp}\|^{2}+4\eta^{2}L_{f}^{2}\|\mathbf{\overline{x}}_{\parallel}\|^{2}\big{)},$
for $c_{4}:=NL_{\phi}\theta$, where the first inequality is obtained (60) and (61), and $r>0$ is a parameter which will be determined later.
We define the same $V_{t}$ as that in Appendix .8, then (59) holds and we have
${rcl}\mathbb{E}\Delta V_{t}\leq-\gamma V_{t},\ \gamma=\frac{1}{2}\kappa\mathrm{min}\{\frac{\lambda_{2}}{2},{\beta^{2}},\eta\frac{\mu_{n}}{2},\frac{c_{3}}{2c_{1}}\}.$
with the same parameters. Then we can derive $\mathbb{E}\|\overline{\mathbf{x}}(t)\|^{2}=\mathcal{O}((1-\gamma)^{t})$. With the definition $\overline{\mathbf{x}}=\mathbf{x}-\mathbf{s}$ before, we know that the mean square of $\mathbf{x}_{i}(t)$ in DPD-OC converge exponentially to the optimal solution $s^{\ast}$ with the StST compressor.